### v1

# GenIAS: Generator for Instantiating Anomalies in time Series

Abstract.
A recent and promising approach for building time series anomaly detection (TSAD) models is to inject synthetic samples of anomalies within real data sets. The existing injection mechanisms have significant limitations — most of them rely on ad hoc, hand-crafted strategies which fail to capture the natural diversity of anomalous patterns, or are restricted to univariate time series settings. To address these challenges, we design a generative model for TSAD using a variational autoencoder, which is referred to as a Generator for Instantiating Anomalies in Time Series (GenIAS). GenIAS is designed to produce diverse and realistic synthetic anomalies for TSAD tasks. By employing a novel learned perturbation mechanism in the latent space and injecting the perturbed patterns in different segments of time series, GenIAS can generate anomalies with greater diversity and varying scales. Further, guided by a new triplet loss function, which uses a min-max margin and a new variance-scaling approach to further enforce the learning of compact normal patterns, GenIAS ensures that anomalies are distinct from normal samples while remaining realistic. The approach is effective for both univariate and multivariate time series. We demonstrate the diversity and realism of the generated anomalies. Our extensive experiments demonstrate that GenIAS — when integrated into a TSAD task — consistently outperforms seventeen traditional and deep anomaly detection models, thereby highlighting the potential of generative models for time series anomaly generation.
Anomaly generator, Time series, Anomaly detection
1. Introduction
Unsupervised methods for Time Series Anomaly Detection (TSAD) often learn a model for boundary of normal data without access to labeled anomalies. This reliance on normal data makes models vulnerable to unseen or diverse anomaly patterns, which are common in time series settings. This is particularly true in complex multivariate scenarios, where anomalous patterns may vary significantly across different dimensions or even be correlated.
Anomaly injection techniques are an increasingly popular approach to address the limitations of solely relying on normal-boundary learning. Models like CARLA (Darban et al., 2025), CutAddPaste (Wang et al., 2024), NCAD (Carmona et al., 2022), and COUTA (Xu et al., 2024) have demonstrated the potential of synthetic anomaly injection in improving detection capabilities. However, these models face notable limitations. CARLA, CutAddPaste, and COUTA are restricted to predefined anomaly types, such as point, seasonal, trend, and shapelet anomalies. CutAddPaste further employs a rigid augmentation strategy (of “cut-add-paste”), which may not align well with realistic patterns in complex applications. NCAD, on the other hand, relies on prior knowledge about anomalies to generate realistic anomalies. However, it is limited by the anomalies present in the datasets and hence may not be able to generate diverse types of anomalies. Finally, all the existing anomaly injection models are susceptible to slight variations in the time series, as they generate anomalies in the raw time series space.
We address these gaps by introducing the Generator for Instantiating Anomalies in Time Series (GenIAS), a novel generative model designed to produce diverse and realistic synthetic anomalies for TSAD. In contrast to previous approaches, GenIAS operates in the latent space rather than the raw time series space. This makes it less susceptible to minor variations in the raw time series and allows it to learn more generalized normal patterns. While generative models mainly used to generate normal data, we are able to leverage the power of variational neural models to generate anomalies that align well with the temporal properties of time series, and with correlations between variables if the input time series is multivariate. GenIAS builds upon a hybrid Temporal Convolutional Network-Variational Autoencoder (TCN-VAE) architecture to move beyond rigid, handcrafted augmentation strategies and maps the complex properties of time series data into a multidimensional latent space. By employing a learned perturbation mechanism applied to the latent space variance and injecting the perturbed patterns into different segments of the time series using a patching mechanism, GenIAS generates anomalies with varying scales and enhanced diversity, resulting in more realistic anomaly patterns. We use the term realism to refer to realistic anomaly patterns in this paper. Additionally, we employ an adjusted triplet loss function, guided by a min-max margin to guarantee that the generated anomalies are both distinct from normal data and realistic at the same time. Further, we introduce a novel variance-scaling approach, which enforces the compactness of latent representations, making the generated anomalies more distinguishable from normal samples. We demonstrate the effectiveness of GenIAS when extended as a TSAD model. Our key contributions are as follows:
  • We propose GenIAS, which is a novel anomaly generator designed to seamlessly handle both univariate time series (UTS) and multivariate time series (MTS). By employing deep generative techniques, GenIAS produces anomalies that are both diverse and realistic.
  • We propose a novel loss function for our VAE-based anomaly generator, enforcing compact latent representations via a tailored prior variance. This improves the separation between normal and anomalous data in latent space, enhancing anomaly generation (see Figure 1). A theoretical analysis supports its effectiveness.
  • We demonstrate GenIAS’s ability to generate diverse and realistic anomalies through two newly introduced metrics. The evaluation is conducted on four state-of-the-art TSAD models leveraging anomaly injection, across nine widely used MTS and UTS datasets.
  • We demonstrate GenIAS’s broad applicability by integrating it into a state-of-the-art TSAD model and comparing it with 17 baselines. Extensive experiments on nine diverse UTS and MTS datasets confirm its effectiveness, with performance gains directly linked to the quality of GenIAS’s anomaly generation.
Figure 1. An illustration of our key idea: Variance-scaling enforces latent space compactness in the latent space, strengthening the separation between normal and generated anomalous samples by GenIAS, in latent space. The bottom figure, with $\sigma_{\text{prior}}$ = 0.5, visually shows this effect.
2. Related Work
The field of time series anomaly detection has seen remarkable progress, and include traditional statistical methods as well as state-of-the-art deep learning approaches (Schmidl et al., 2022; Audibert et al., 2022; Zamanzadeh Darban et al., 2024). Early efforts included models like One-Class SVM (OC-SVM) (Schölkopf et al., 1999), LOF (Breunig et al., 2000), and Isolation Forest (Fei Tony Liu and Zhou, 2008); however, these methods faced challenges in handling high-dimensional and sequential data. Deep learning models like Donut (Xu et al., 2018), LSTM-VAE (Park et al., 2018) and OmniAnomaly (Su et al., 2019) introduced temporal dependency modeling and probabilistic approaches in TSAD.
Representation-based methods, have pushed TSAD forward by focusing on meaningful embeddings. TS2Vec (Yue et al., 2022) applies self-supervised learning to capture multi-level semantic representations, while DCdetector (Yang et al., 2023) employs dual attention for permutation-invariant representations. Recent models such as TimesNet (Wu et al., 2023) and TranAD (Tuli et al., 2022a) model complex dependencies by leveraging attention mechanisms. Methods like MTAD-GAT (Zhao et al., 2020) utilize graph attention networks to model multivariate dependencies and THOC (Shen et al., 2020a) incorporates hierarchical modeling to capture temporal patterns at multiple granularities. Anomaly Transformer (Xu et al., 2021) leverages anomaly-sensitive features through attention mechanisms.
OE (Hendrycks et al., 2019) integrates external data during training to expose the model to out-of-distribution instances, thereby making it more powerful. NCAD(Carmona et al., 2022) adapts OE for time series by infusing mixed contextual and random point anomalies to better separate normal and anomalous patterns. Anomaly injection and perturbation methods have emerged as crucial strategies for enhancing TSAD models. Approaches like CutAddPaste (Wang et al., 2024) and COUTA (Xu et al., 2024) simulate diverse anomaly patterns during training. These techniques have also been integrated into advanced models like CARLA (Darban et al., 2025) to create negative pairs for contrastive learning, enhancing generalization to unseen anomaly types.
3. GenIAS Method
3.1. Problem Statement
Let $\mathcal{D}$ represent a time series dataset consisting of $N$ time series windows, where each window $\mathbf{X}\in\mathbb{R}^{T\times D}$ is a time series window of length $T$ with $D$ dimension(s). Assume that samples $X$ in $\mathcal{D}$ is drawn from an unknown data distribution $p(\mathbf{X})$. In TSAD datasets, the training data is typically unlabeled, and it is assumed normal since anomalies are rare and have negligible impact. The objective is to use a generative mapping $\Gamma:\mathbf{X}\rightarrow\widetilde{\mathbf{X}}$ to transform a normal sample $\mathbf{X}$ into an anomalous sample $\widetilde{\mathbf{X}}\in\mathbb{R}^{T\times D}$, such that the anomalous distribution $p(\widetilde{\mathbf{X}})$ satisfies $p(\widetilde{\mathbf{X}})\neq p(\mathbf{X})$. The anomalies must preserve essential temporal properties in the normal data while exhibiting deviations that make them statistically separated.
The first challenge is to define a perturbation mechanism on the latent representation $\mathbf{z}\in\mathbb{R}^{L}$ with $L$ dimensions, parameterized by $p(\mathbf{z}\mid\mathbf{X})$, to produce a shifted latent distribution $p(\widetilde{\mathbf{z}})$. This ensures that the anomalies are distinguishable, satisfying $p(\widetilde{\mathbf{z}})\neq p(\mathbf{z})$ while retaining plausibility by remaining within the data manifold. Furthermore, by perturbing the latent representations, our model is less susceptible to the slight deviations (e.g., noise) in the raw time series space. The second challenge is to densify the normal latent distribution $p(\mathbf{z})$, enabling the mapping $\Gamma(\mathbf{X})=\hat{\mathbf{X}}$ to accurately reconstruct normal samples and establish a clear boundary for generating anomalies. Together, addressing these challenges guarantees that the generated anomalies achieve a balance between diversity and realism. This approach is designed to handle univariate ($D=1$) and multivariate ($D>1$) time series.
3.2. Anomaly Generation
GenIAS is a generative model designed to learn compact representations of normal time series data and generate realistic anomalies by strategically perturbing the latent space. The model comprises three main components: an encoder, a latent space representation, and a decoder. It is trained end-to-end with an objective function that balances reconstruction accuracy, latent space regularization, and anomaly generation quality.
3.2.1. Model Workflow and Perturbation Mechanism
The input to the model is a time series window $\mathbf{X}$ of shape $T\times D$. The input $\mathbf{X}$ is normalized to ensure consistency across features, improving model stability during training. The model architecture combines a TCN and a VAE to combine the strengths of both approaches. The VAE learns a latent representation of normal data and perturbs it in the latent space to enhance robustness to noisy time series and improve the learning of normal patterns. The TCN captures temporal dependencies within the time series more effectively. The encoder maps the input sample $\mathbf{X}$ into a latent representation $\mathbf{z}$ by capturing temporal dependencies through stacked convolutional layers. These layers, interleaved with activation functions and dropout, effectively model complex temporal patterns while reducing overfitting. The encoder outputs the parameters $\mu$ (mean) and $\sigma^{2}$ (variance), which define a Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$ in the latent space, where $\mu$ and $\sigma$ are vectors of size $L$, the same as the size of latent vector $\mathbf{z}$.
The latent space representation forms the core of the model. For normal samples, latent vectors $\mathbf{z}$ are sampled as: $\mathbf{z}\sim\mathcal{N}(\mu,\sigma^{2})$. To generate anomalous representations $\tilde{\mathbf{z}}$, the standard deviation $\sigma$ is perturbed while keeping $\mu$ unchanged:
(1) $\tilde{\mathbf{z}}=\mu+\psi\cdot(\sigma\odot\epsilon),$
where $\psi$ is the learned perturbation scale, and $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is a noise vector. Perturbing $\sigma$ instead of $\mu$ ensures that the anomalies deviate in a way that preserves the central tendency of the normal data, thereby maintaining realism. Altering $\mu$ could disrupt this central tendency, leading to unrealistic deviations that do not align with the underlying data manifold.
The decoder reconstructs the input sample from $\mathbf{z}$ and $\tilde{\mathbf{z}}$. Using transpose convolutional layers upscales the latent representations back to the original input dimensions while preserving the temporal structure. The decoder outputs the following:
  • $\hat{\mathbf{X}}$: The reconstructed normal window, derived from $\mathbf{z}$.
  • $\tilde{\mathbf{X}}$: The generated anomalous window, derived from $\tilde{\mathbf{z}}$.
Figure 2 illustrates the overall workflow of the model, including the TCN-VAE architecture and the perturbation mechanism. The training process optimizes a combination of losses to ensure accurate reconstruction, realistic anomaly generation, and a compact latent space learning (see Algorithm 1 for details).
Figure 2. The overall architecture of GenIAS: The model comprises an encoder, a structured latent space, and a decoder. The encoder, built with a TCN-VAE, maps input time series windows into a latent representation by learning temporal dependencies. The latent space regularization ensures compact normal representations, while anomalies are generated by perturbing the variance of the learned distribution without altering the mean, preserving realistic deviations. The decoder reconstructs both normal and perturbed windows using transpose convolutional layers.
3.2.2. Reconstruction Loss
This portion of the overall loss is designed to accurately reconstruct normal input samples. It is defined using the Mean Squared Error (MSE) between an input sample $\mathbf{X}$ and its reconstruction $\hat{\mathbf{X}}$:
(2) $\mathcal{L}_{\text{recon}}=\frac{1}{|\mathcal{D}|}\sum_{\mathbf{X}\in\mathcal{D}}\frac{1}{T\cdot D}\sum_{t=1}^{T}\sum_{d=1}^{D}\big{(}\mathbf{X}[t,d]-\hat{\mathbf{X}}[t,d]\big{)}^{2},$
This loss penalizes the discrepancy between $\mathbf{X}$ and $\hat{\mathbf{X}}$, ensuring that the model captures the key temporal and structural features of normal samples.
3.2.3. Perturbation Loss
The perturbation loss ensures that the generated anomalies $\tilde{\mathbf{X}}$ are both distinct from normal samples $\mathbf{X}$ and exhibit realistic deviations. This is achieved through two complementary terms: a triplet loss to enforce distinctiveness and a regularization term to maintain realism.
The first term, the triplet loss, encourages the original time series sample ${\mathbf{X}}$ to be closer to the reconstructed sample $\hat{\mathbf{X}}$ than the generated anomalous sample $\tilde{\mathbf{X}}$, ensuring that the anomalies are sufficiently distinct from original time series. By maximizing the margin between $d(\mathbf{X},\hat{\mathbf{X}})$ and $d(\mathbf{X},\tilde{\mathbf{X}})$, where $d(x,y)$ denotes the MSE distance, this term pushes the anomalous samples further away from the normal samples in the raw time series space. We also ensure that this margin is at least $\delta_{\text{min}}$, thus the generated anomalies be more distinct.
The second term (the regularization term) prevents the anomalous samples $\tilde{\mathbf{X}}$ from deviating excessively from the original samples, thereby preserving realism. It penalizes deviations where distance $d(\mathbf{X},\tilde{\mathbf{X}})$ exceeds a maximum allowable threshold $\delta_{\text{max}}$. This ensures that the generated anomalies remain plausible and do not degenerate into unstructured or unrealistic noise.
The perturbation loss is formulated as:
(3) $\begin{split}\mathcal{L}_{\text{perturb}}=&\frac{1}{|\mathcal{D}|}\sum_{\mathbf{X}\in\mathcal{D}}\Big{[}\max\big{(}d(\mathbf{X},\hat{\mathbf{X}})-d(\mathbf{X},\tilde{\mathbf{X}})+\delta_{\text{min}},0\big{)}\Big{]}\\ &+\frac{1}{|\mathcal{D}|}\sum_{\mathbf{X}\in\mathcal{D}}\max\big{(}d(\mathbf{X},\tilde{\mathbf{X}})-\delta_{\text{max}},0\big{)},\end{split}$
Together, these terms ensure that the generated anomalies achieve the dual objectives of being distinct yet realistic.
3.2.4. Zero-Perturbation Loss
The zero-perturbation loss focuses on dimensions in the input sample $\mathbf{X}$ that are entirely zero. It ensures that anomalies $\tilde{\mathbf{X}}$ introduce meaningful perturbations in these dimensions, encouraging deviations from the original sample where appropriate. The loss is defined as:
(4) $\mathcal{L}_{\text{zero-perturb}}=\frac{1}{|\mathcal{D}|}\sum_{\mathbf{X}\in\mathcal{D}}(\frac{1}{T\cdot|\mathcal{Z}|}\sum_{t=1}^{T}\sum_{d\in\mathcal{Z}}\big{(}d(\mathbf{X},\tilde{\mathbf{X}})+1)^{-1},$
where $\mathcal{Z}$ is the set of zero-valued dimensions in $\mathbf{X}$. By penalizing dimensions that deviate unrealistically in regions of zero value, this loss encourages the model to generate plausible anomalies while preserving the unique characteristics of zero-valued dimensions.
3.2.5. Enhanced KL Divergence Loss
The standard KL divergence loss is typically used to regularize the latent space by aligning the approximate posterior distribution $q(z|\mathbf{X})$, parameterized by $\mu$ and $\sigma^{2}$, with a standard Gaussian prior $p(z)$, defined as $\mathcal{N}(0,1)$. While effective for many tasks, this standard approach may fail to create the tight and compact latent representations necessary for generating anomalies that are distinguishable from normal samples.
To address this, we propose an enhanced KL divergence loss by incorporating a tunable prior variance, $\sigma_{\text{prior}}^{2}$, where $\sigma_{\text{prior}}<1$. This modification enforces tighter representations of normal samples in the latent space. The modified KL divergence loss is defined as:
(5) $\mathcal{L}_{\text{en-KL}}=-\frac{1}{2}\sum_{i=1}^{|\mathbf{\mathcal{D}}|}\sum_{j=1}^{L}\Big{[}1+\log(\sigma^{2}_{ij})-\mu_{ij}^{2}-\frac{\sigma^{2}_{ij}}{\sigma_{\text{prior}}^{2}}+2\log(\sigma_{\text{prior}})\Big{]},$
Where $\mu_{ij}$ and $\sigma^{2}_{ij}$ are the mean and variance of the latent distribution for the $j$-th dimension of the $i$-th sample, and $\sigma_{\text{prior}}$ is the prior’s standard deviation.
Theorem 1 (Enhanced Separation Between Normals and Anomalies in Latent Space).
Let a VAE be trained on normal data with a latent prior $\mathcal{N}(0,\sigma^{2}_{\text{prior}})$, where $\sigma_{\text{prior}}<1$. The encoder posterior for normal samples follows $\mathcal{N}(\mu,\sigma^{2}_{\text{normal}})$, with compactness enforced by KL regularization. If the encoder $\phi$ applies a perturbation scale $\psi$, inflating variance as $\sigma^{2}_{\text{anom}}=\psi\sigma^{2}_{\text{normal}},\quad\text{with }\psi>1,$ then the KL divergence between normal and anomalous latent distributions strictly increases compared to the case where $\sigma_{\text{prior}}\geq 1$.
The proof is provided in Appendix D.3. We also illustrate in Figure 8 in the Appendix how our enhanced KL loss leads to the separation of reconstruction error in the raw space.
The benefits of the enhanced KL divergence loss are as follows:
  • Controlled spread: $\sigma_{\text{prior}}$ determines the spread of normal samples in the latent space. Smaller values enforce tighter clustering, leading to compact representations while allowing limited diversity for variability. (See Lemma 1)
  • Regularization: The KL loss with $\sigma_{\text{prior}}$ ensures a well-regularized latent space, reducing overfitting and promoting smooth sampling.
  • Improved normal sample representation: By constraining the latent space, normal samples occupy a distinct, dense region, enhancing their separability from anomalies. (See Lemma 2 and Theorem 1)
  • Enhanced interpretability: The structured latent space shaped by $\sigma_{\text{prior}}$ provides consistent and interpretable embeddings for both normal and anomalous data.
3.2.6. Total Loss
The total loss integrates all the above components, each weighted by a corresponding hyperparameter:
(6) $\mathcal{L}_{\text{total}}=\alpha\cdot\mathcal{L}_{\text{recon}}+\beta\cdot\mathcal{L}_{\text{perturb}}+\gamma\cdot\mathcal{L}_{\text{zero-perturb}}+\zeta\cdot\mathcal{L}_{\text{en-KL}},$
Where $\alpha$, $\beta$, $\gamma$, and $\zeta$ are the weights for the reconstruction, perturbation, zero-perturbation, and KL divergence losses, respectively.
3.3. Deviation Patching Approach
After training the GenIAS model in the first step, the anomaly injection process uses the pretrained model to perturb the latent space representation of time series windows ($\mathbf{X}$) and generate realistic anomalous samples ($\tilde{\mathbf{X}}$). This step consists of two phases: (1) anomaly injection via latent space perturbation and (2) patching the injected anomalies using a deviation-based approach, defined by a scaling factor $\tau$, to ensure the generated samples are realistic and aligned with the domain requirements. The injection and patching process is applied independently for each dimension ($d$) of the MTS, and more details are provided in Algorithm 2.
Figure 3. Effect of deviation-based patching with different thresholds (0.4, 0.2, 0.05) on generated anomalous windows across various entities from the MSL, SMAP, SMD, Yahoo, and KPI datasets.
4. Experiments
We conduct extensive evaluations of GenIAS’s performance using a comprehensive selection of datasets and baselines. Our evaluation consists of two parts: (1) Generation Evaluation, where we compare GenIAS against other anomaly injection/generation models to assess its capability in generating realistic and diverse anomalies, and (2) Case Study – TSAD Evaluation, where we benchmark GenIAS against seventeen competing methods using nine real-world datasets covering both MTS and UTS. Specifically, to study the effectiveness of GenIAS’s anomaly generation in TSAD, we integrate GenIAS with the anomaly detector architecture from CARLA to implement a GenIAS-enhanced TSAD model (i.e., substituting CARLA’s anomaly generator with GenIAS while keeping the rest of the framework identical). We choose CARLA as our base detector due to its SOTA performance, which enables direct comparison with the strongest generation-based TSAD method. Through this comprehensive evaluation, we provide a thorough analysis of GenIAS’s capabilities in both anomaly generation and TSAD tasks.
Table 1. An overview of dataset statistics.
Benchmark # datasets # dims Train size Test size ARMSL (Hundman et al., 2018) 27 55 58,317 73,729 10.72% SMAP (Hundman et al., 2018) 55 25 140,825 444,035 13.13% SMD (Su et al., 2019) 28 38 708,405 708,420 4.16% SWaT (Mathur and Tippenhauer, 2016) 1 51 496,800 449,919 12.33% GECCO (Moritz et al., 2018) 1 9 69,260 69,261 1.1% SWAN (Angryk et al., 2020) 1 38 60,000 60,000 32.6% UCR (Wu and Keogh, 2021) 205 1 2,238,349 6,143,541 0.6% Yahoo-A1 (Laptev et al., 2015) 67 1 46,667 46,717 1.76% KPI (aio, 2018) 29 1 1,048,576 2,918,847 1.87%
4.1. Benchmark Datasets
GenIAS is thoroughly evaluated on nine widely adopted real-world benchmark datasets for TSAD. Their key statistics are reported in Table 1. In particular, our selection includes six MTS datasets (MSL, SMAP, SMD, SWaT, GECCO, and SWAN) and three UTS datasets (UCR, Yahool-A1 and KPI), containing diverse anomaly types and a variety of anomaly ratios, thereby providing a comprehensive evaluation of the performance of our proposed methods and the baselines. Detailed dataset descriptions can be found in Appendix A.
4.2. Baselines
GenIAS is compared with seventeen relevant and competitive baseline methods that leverage a wide array of anomaly detection approaches, which can be categorised into four broad categories: classic anomaly detection methods, positive-unlabeled (P-U) methods, unsupervised deep TSAD methods that do not leverage anomaly generation (US-WI), and unsupervised TSAD methods that employ anomaly generation (US-I).
Specifically, three classic general AD methods - OCSVM (Schölkopf et al., 1999), LOF (Breunig et al., 2000), and Isolation Forest (Fei Tony Liu and Zhou, 2008) - are included to establish a base performance for demonstrating the advantages of deep learning TSAD approaches. LSTM-VAE (Park et al., 2018), a deep learning TSAD method that utilises labeled anomalies, is included to examine the effects of leveraging labelled anomalies for TSAD. US-WI is another relevant category, which includes a variety of unsupervised TSAD baselines that do not leverage anomaly generation. Donut (Xu et al., 2018), OmniAnomaly (Su et al., 2019), AnomalyTransformer (Xu et al., 2021), and TranAD (Tuli et al., 2022b) use reconstruction error to detect anomalies. THOC (Shen et al., 2020b) and TimesNet (Wu et al., 2023) treat TSAD as a forecasting problem. MTAD-GAT (Zhao et al., 2020) integrates graph attention mechanisms to capture both temporal and multivariate dependencies. Representation learning methods like TS2Vec (Yue et al., 2022) and DCdetector (Yang et al., 2023) focus on establishing robust and anomaly-discriminative representations. US-I is the most relevant category, which allows us to compare GenIAS with existing generation methods for unsupervised TSAD. We include four recent unsupervised TSAD methods that leverage various anomaly generation approaches: NCAD (Carmona et al., 2022), CutAddPaste (Wang et al., 2024), COUTA (Xu et al., 2024), and CARLA (Darban et al., 2025).
4.3. Evaluation Metrics
We adopt seven evaluation metrics to assess GenIAS’s performance in terms of generation quality and TSAD performance.
Generation quality. To quantitatively examine the quality of anomaly generation, we introduce two key metrics: Anomalous Representation Proximity (ARP), which assesses realism, and Entropy-Based Diversity Index (EDI), which characterizes the diversity of the generated anomalies, motivated by the idea in AnomalyDiffusion (Hu et al., 2024). Due to page limits, the definitions and discussions of these metrics are provided in the Appendix E. In general, ARP is the primary metric as it measures the closeness of the generated anomalies to the actual anomalies, see Appendix E.1. EDI is complementary to ARP, as diversity should only be examined under the condition that the generated anomalies can illustrate the actual anomalies, see Appendix E.2.
TSAD evaluation. We consider five widely-used, complementary metrics: Best F1 score, AUPR (Area Under the Precision-Recall Curve), Affiliation F1  (Huet et al., 2022), PATE  (Ghorbani et al., 2024), and AUROC (Area Under the Receiver Operating Characteristic Curve) for TSAD performance evaluation. This combination of metrics provides a comprehensive, fair and robust TSAD evaluation. Detailed descriptions of evaluation metrics are provided in Appendix E.3.
4.4. Implementation Details
Our experiments are conducted on a server with an NVIDIA A40 GPU, 13 CPU cores, and 250GB RAM. We employ the unsupervised TSAD network architecture and training scheme from CARLA as our default TSAD model to perform GenIAS-enhanced detection. GenIAS is implemented in Python using PyTorch and trained using the Adam optimizer with a learning rate of $10^{-4}$ for 1000 epochs. The default window size $T$ is set to 200, and latent dimension size $L$ is set to 100 and 50, for MTS and UTS respectively. The hyperparameters for the total loss are $\alpha=1.0$, $\beta=0.1$, $\gamma=0.0/0.01$ (for UTS/MTS), and $\zeta=0.1$, which are fixed for all datasets. We adopt the baseline results that are produced using the original implementation and maintain their default hyperparameter settings. Due to the page limit, the detailed implementation details and hyperparameter settings are provided in Appendix F.
Table 2. Anomaly generation quality of GenIAS and US-I baselines in ARP, with the best model in bold and the second-best model underlined, where CAP abbreviates CutAddPaste.
Metric Dataset COUTA NCAD CAP CARLA GenIAS ARP MSL 0.3613 0.3568 0.3406 0.3586 0.3605 SMAP 0.7573 0.8052 0.8110 0.7636 0.8113 SMD 0.0951 0.0810 0.0896 0.0958 0.0922 SWaT 0.9970 0.9964 0.9963 0.9980 0.9974 GECCO 0.1296 0.1145 0.1335 0.1163 0.1265 SWAN 0.2400 0.2362 0.2313 0.2410 0.2415 Yahoo 0.6541 0.6139 0.6476 0.6632 0.6756 KPI 0.6856 0.8427 0.7207 0.7582 0.7907 UCR 0.1998 0.2059 0.1936 0.2086 0.2161 Rank 3.11 3.78 3.89 2.44 1.78 EDI MSL 0.8251 0.7311 0.7169 0.5822 0.8833 SMAP 0.8557 0.6699 0.6364 0.6335 0.8932 SMD 0.8782 0.7648 0.6404 0.8041 0.9262 SWaT 0.6425 0.6164 0.7967 0.6005 0.7259 GECCO 0.7379 0.5571 0.7072 0.9327 0.8303 SWAN 0.9589 0.9189 0.7428 0.9286 0.9840 Yahoo 0.8627 0.7202 0.7577 0.5948 0.9182 KPI 0.7607 0.6166 0.7505 0.6962 0.8442 UCR 0.8166 0.7321 0.7536 0.7483 0.8961 Rank 2.22 4.11 3.56 3.89 1.22
4.5. Generation Performance
We first examine the anomaly generation quality of GenIAS and the competing methods that employ anomaly injection (i.e., the US-I category, see section 4.2) in terms of ARP and EDI, as discussed in Section 4.3. We report the results of ARP and EDI in Table 2. GenIAS achieves the highest average ARP ranking, suggesting it generates overall the most realistic anomalies across all datasets. The realism of the anomaly distribution is essential for providing anomaly-informed supervision in the absence of labelled anomalies. Similarly for EDI, GenIAS also achieves the highest average ranking, indicating good diversity in its generated anomalies while sufficiently illustrating the actual anomalies. This diversity implies that the generated anomalies can also lie beyond the original anomaly distribution, which can be beneficial for learning a more stringent normal class boundary, further improving TSAD performance, provided that they do not interfere with the identification of the normal region. To achieve this, GenIAS treats creating informative but not excessive diversity as a key consideration through its novel perturbation loss.
Table 3. TSAD performance of GenIAS and baseline methods using F1 score and AUPR, with the best model in bold and the second-best model underlined. U and M denote results unavailable for models limited to UTS or MTS, respectively.
MTS UTS Metric Cat. Methods MSL SMAP SMD SWAT GECCO SWAN Avg-M Yahoo KPI UCR Avg-U Avg F1 Classic OCSVM (1999) 0.308 0.411 0.319 0.743 0.297 0.557 0.439 0.608 0.148 0.011 0.256 0.378 LOF (2000) 0.368 0.400 0.204 0.331 0.075 0.655 0.339 0.586 0.101 0.010 0.232 0.303 IForest (2008) 0.278 0.353 0.306 0.745 0.255 0.717 0.442 0.553 0.236 0.010 0.266 0.384 PU LSTM-VAE (2018) 0.407 0.437 0.298 0.724 0.208 0.625 0.450 M M M M M US-WI Donut (2018) U U U U U U U 0.489 0.126 0.009 0.208 U Omni (2019) 0.243 0.325 0.459 0.763 0.166 0.541 0.416 M M M M M THOC (2020) 0.309 0.327 0.168 0.638 0.146 0.506 0.349 0.253 0.233 0.010 0.165 0.288 MTAD-GAT(2020) 0.473 0.519 0.347 0.242 0.313 0.599 0.416 M M M M M AnomalyTran (2021) 0.345 0.407 0.304 0.738 0.155 0.524 0.412 M M M M M TranAD (2022) 0.428 0.472 0.361 0.310 0.280 0.527 0.396 0.566 0.287 0.010 0.288 0.360 TS2Vec (2022) 0.299 0.371 0.173 0.261 0.063 0.502 0.278 0.484 0.204 0.007 0.232 0.263 TimesNet (2023) 0.358 0.402 0.338 0.217 0.297 0.492 0.351 0.513 0.241 0.011 0.255 0.319 Dcdetector (2023) 0.227 0.275 0.083 0.217 0.021 0.492 0.219 0.112 0.042 0.008 0.054 0.164 US-I NCAD (2022) 0.266 0.358 0.183 0.217 0.296 0.492 0.302 0.134 0.168 0.020 0.107 0.237 CutAddPaste (2024) 0.363 0.458 0.192 0.700 0.280 0.896 0.482 0.321 0.620 0.042 0.328 0.430 COUTA (2024) 0.414 0.421 0.362 0.312 0.389 0.626 0.421 0.658 0.221 0.010 0.296 0.379 CARLA (2025) 0.523 0.529 0.511 0.721 0.293 0.908 0.581 0.749 0.308 0.039 0.365 0.509 Ours GenIAS 0.579 0.620 0.528 0.755 0.329 0.937 0.625 0.811 0.393 0.064 0.423 0.557 AUPR Classic OCSVM (1999) 0.202 0.256 0.327 0.696 0.163 0.608 0.375 0.703 0.300 0.028 0.344 0.365 LOF (2000) 0.217 0.248 0.277 0.185 0.031 0.624 0.264 0.615 0.109 0.043 0.256 0.261 IForest (2008) 0.182 0.190 0.278 0.739 0.162 0.739 0.382 0.614 0.288 0.021 0.308 0.357 PU LSTM-VAE (2018) 0.285 0.258 0.395 0.685 0.045 0.674 0.390 M M M M M US-WI Donut (2018) U U U U U U U 0.264 0.078 0.015 0.119 U Omni (2019) 0.149 0.115 0.365 0.713 0.034 0.587 0.327 M M M M M THOC (2020) 0.240 0.195 0.107 0.537 0.076 0.432 0.265 0.349 0.229 0.014 0.197 0.242 MTAD-GAT(2020) 0.335 0.339 0.401 0.095 0.200 0.613 0.331 M M M M M AnomalyTran (2021) 0.236 0.264 0.273 0.680 0.059 0.568 0.347 M M M M M TranAD (2022) 0.273 0.287 0.412 0.192 0.119 0.586 0.312 0.691 0.285 0.022 0.333 0.319 TS2Vec (2022) 0.132 0.148 0.113 0.136 0.040 0.408 0.163 0.491 0.221 0.009 0.240 0.189 TimesNet (2023) 0.283 0.208 0.385 0.082 0.189 0.508 0.276 0.671 0.237 0.032 0.313 0.288 Dcdetector (2023) 0.129 0.124 0.043 0.126 0.011 0.326 0.127 0.041 0.018 0.009 0.023 0.092 US-I NCAD (2022) 0.146 0.156 0.096 0.106 0.135 0.407 0.174 0.067 0.089 0.103 0.086 0.145 CutAddPaste (2024) 0.242 0.322 0.153 0.582 0.159 0.808 0.378 0.116 0.602 0.188 0.302 0.352 COUTA (2024) 0.247 0.248 0.400 0.159 0.264 0.685 0.334 0.555 0.282 0.049 0.295 0.321 CARLA (2025) 0.501 0.448 0.507 0.681 0.201 0.814 0.525 0.627 0.299 0.247 0.391 0.481 Ours GenIAS 0.529 0.467 0.512 0.704 0.236 0.882 0.555 0.794 0.357 0.280 0.477 0.529
4.6. Visualization
(a) Yahoo - Entity real_65
(b) SMAP - Entity E5
Figure 4. Each plot visualizes a different time series (TS) injection or generation method using t-SNE. Blue markers represent the original anomalous TS, while orange markers represent the injected or generated anomalous TS.
To illustrate GenIAS’s enhanced anomaly generation compared to baseline methods that also leverage anomaly generation (US-I), we present t-SNE visualizations comparing their generated and real anomalies in Figure 4 on two selected datasets, Yahoo (UTS) and SMAP (MTS). Some additional visualizations are included in Figure 10 in Appendix G.2.3.
Impressively, the distributions of the generated anomalies by GenIAS align with the actual anomalies distribution substantially closer compared to the baseline methods (i.e., actual anomaly regions are densely covered by the generated anomalies). This is essential to provide anomaly-discriminative information for establishing more distinctive normal and anomaly representations in unsupervised TSAD training. On the other hand, the baseline methods, to varying degrees, are unable to generate anomalies that sufficiently represent the actual anomaly distribution, which is evidenced by major blue areas (representing actual anomaly regions) that are not covered by orange areas (generated anomaly regions). For example, CARLA, the second-best method for overall ARP and TSAD performance, also suffers from these issues: on Yahoo (a), generated anomalies largely miss the anomaly regions, while on SMAP (b), generated anomalies form two major clusters leaving a gap in between, which is an actual anomaly region failing to be represented. These observations further substantiate that relying on handcrafted, predefined anomaly patterns for anomaly generation is prone to distribution misalignment when the assumed patterns deviate from actual ones, and disconnected/scattered anomaly regions when the assumed patterns differ significantly. Whereas GenIAS, through controlled and regularized perturbation on the learned distribution, generates anomalies without pattern speculation, achieving both better alignment and smoothness.
4.7. Main TSAD Performance
We report the TSAD performance of GenIAS and the competing methods in terms of the best F1 score and AUPR in Table 3. Due to space limitations, please refer to the Appendix G.2.2 for results on the other three previously mentioned metrics discussed in Section 4.3. Impressively, GenIAS achieves significantly better overall TSAD performance compared to the baseline methods across all metrics, demonstrating its superior effectiveness in handling both MTS and UTS. Specifically, among methods that can operate on both types of time series, GenIAS achieves 9.4% and 29.5% performance improvements compared to the second and third best performing methods in F1, respectively. Similarly, it outperforms the second best model by 10.0% and third best model by 44.9% in terms of AUPR. Upon examining the more detailed TSAD performance with respect to factors such as method category and dataset type, we noted the following observations.
Consistent effectiveness for both MTS and UTS. On each type of dataset, GenIAS yields similar levels of performance improvement as observed in overall performance. For example, compared to the second-best model for both types, GenIAS improves F1 by 7.6% and AUPR by 8.4% on MTS, and improves F1 by 15.9% and AUPR by 22.0% for UTS. This highlights its ability to generate informative and diverse anomalies that can mimic anomalous patterns within each dimension and the dependencies between dimensions. We also provide the average rank of all models in the form of critical difference diagrams (Demšar, 2006) for F1 and AUPR in both MTS and UTS (see Appendix G.2.1). GenIAS achieved the highest rank in all four comparisons and demonstrated statistically significant improvements over all baselines in MTS for both AUPR and F1.
Enhanced Anomaly Generation. It is worth noting that even when compared to US-I baselines, GenIAS consistently outperforms them due to its ability to generate more realistic and diverse anomalous patterns. In particular, since GenIAS-enhanced TSAD model and CARLA only differ in their anomaly generation while sharing the same architecture and training methods, their performance gap underscores the effectiveness of GenIAS’s approach. This is attributed to its TSAD tailored variational representation learning and novel perturbation strategies, which enforce more distinctive normal representations and are not confined to predefined, handcrafted anomaly prototypes.
Importance of anomaly generation in TSAD. It is notable that US-I injection methods are generally more effective than methods in other categories that do not leverage labeled anomalies. Remarkably, some of these methods, such as CutAddPaste and CARLA, can even outperform the supervised baseline LASM-VAE. This suggests that, with well-designed anomaly injection techniques, the generated pseudo anomalies can be sufficiently realistic to represent actual anomalies, including those not well captured by the labeled training data in supervised settings. Moreover, leveraging anomaly generation can generate both larger numbers and more diverse types of anomalies, which can improve TSAD representation learning. Therefore, anomaly injection should be incorporated as an essential module in TSAD models.
4.8. Ablation Study
To better understand the contributions of different components and configurations in our proposed model, we conducted an ablation study. Our analysis focuses on the following aspects: (i) The impact of the patch selection, (ii) The effectiveness of the prior variance, and (iii) The effectiveness of the min-max margin.
4.8.1. Impact of the Patch Selection
First, in Figure 5, we present our study of F1 and AUPR for two patching methods—Length-driven and Deviation patching—evaluated across varying portions of patch lengths and thresholds. The x-axis represents the portion of patch length (ranging from 0.2 to 1, in intervals of 0.2), while thresholds $\tau$ are explored at 0.05, 0.1, 0.2, 0.4, and 0.6. The best results for each method, identified through this parameter exploration, are then compared in Table 4. The results in Table 4 demonstrate that Deviation patching consistently outperforms the other approaches, achieving the highest F1 and AUPR scores across all datasets. For example, it achieves a 0.044 increase in F1 and a 0.063 improvement in AUPR for SMAP compared to the baseline without patching (wo Patch). Length-driven patching, which introduces anomalies based on predefined temporal lengths, shows moderate improvements over the baseline in most scenarios but slightly underperforms on more complex datasets, such as Yahoo, where it reduces F1 and AUPR scores. The strong performance of Deviation patching, as defined in Section 3.3, underscores its ability to better align generated anomalies with true anomaly distributions, making it a robust and effective approach for anomaly generation and detection.
Table 4. Performance metrics for different patching methods.
Metric Dataset wo Patch Length-driven Deviation-based F1 MSL 0.535 0.547 0.579 SMAP 0.533 0.571 0.620 Yahoo 0.755 0.741 0.811 AUPR MSL 0.466 0.488 0.529 SMAP 0.425 0.429 0.467 Yahoo 0.776 0.766 0.794
Figure 5. F1 and AUPR for different patching methods.
4.8.2. Effectiveness of the Prior Variance
Figure 6 analyzes the impact of prior variance on anomaly detection performance, showing consistent trends across datasets and indicating that the model is not highly sensitive to this parameter. For the best F1 scores, prior variances of 0.5 achieve optimal performance across all datasets (MSL: 0.579, SMAP: 0.62, Yahoo: 0.811), indicating a spot where the model balances bias and variance effectively. However, as the prior variance increases beyond 0.5, performance declines, suggesting that overly broad priors introduce excessive flexibility, reducing model specificity. Similarly, the AUPR results follow this trend, with 0.5 showing superior means across datasets and narrower standard deviation (std) bars, emphasizing the robustness of this setting. The increasing std with higher variances further supports that larger priors may lead to unstable and less reliable performance. Overall, a prior variance of 0.5 emerges as the most effective choice, yielding a balance between accuracy and stability for both F1 and AUPR.
Figure 6. F1 and AUPR for different prior variance ($\sigma_{\text{prior}}^{2}$)
4.8.3. Effectiveness of the Min-Max Margin
Figure 7. Heatmaps of F1 and AUPR scores across datasets illustrating the effectiveness of min-max margins.
The heatmaps in Figure 7 illustrate the impact of min-max margins ($\delta_{\text{min}},\delta_{\text{max}}$) on performance metrics, such as F1 and AUPR, across MSL, SMAP, and Yahoo datasets. Smaller max margins ($\delta_{\text{max}}=0.2$) and moderate min margins ($\delta_{\text{min}}=0.1$) consistently yield the best results, with peak F1 scores of 0.579, 0.62, and 0.811 for MSL, SMAP, and Yahoo, respectively. Similarly, the AUPR values align with this trend. These results validate the design rationale in Section 3.2.3, where $\delta_{\text{min}}$ enforces distinctness, and $\delta_{\text{max}}$ maintains realism. Larger margins (e.g., $\delta_{\text{max}}=1$) reduce performance, as seen in the lower F1 scores of MSL and Yahoo. These findings emphasize the importance of balancing min-max margins.
5. Conclusion and Future Work
In this work, we introduced GenIAS, a generative anomaly injection framework designed to enhance time series anomaly detection by creating realistic and diverse anomalies. By perturbing latent representations of a generative model trained on normal data, GenIAS produces anomalies that are similar to real-world anomalies. When integrated into a TSAD model, it further enhances anomaly detection performance by adding high-quality synthetic anomalies to the training set. For future work, we aim to refine adaptive perturbation techniques and an adaptive patching approach to better align with real-world domain-specific anomalies. Additionally, we plan to incorporate correlation between dimensions to guarantee the generation of intermetric anomalies, making the synthetic anomalies more representative of real-world multivariate dependencies. These enhancements will improve the realism of generated anomalies by GenIAS and its effectiveness in time series anomaly detection.
Appendix A Datasets Descriptions
Specifically, five popular TSAD datasets are included to evaluate multivariate TSAD. Among these, NASA Datasets — The Mars Science Laboratory (MSL) and Soil Moisture Active Passive (SMAP) datasets [^1] https://www.kaggle.com/datasets/patrickfleith/nasa-anomaly-detection-dataset-smap-msl (Hundman et al., 2018) are collected from NASA spacecraft. These datasets include anomaly information derived from incident reports for spacecraft monitoring systems. Server Machine Dataset (SMD) [^2] https://github.com/NetManAIOps/OmniAnomaly/tree/master/ServerMachineDataset (Su et al., 2019) comprises data from 28 servers over 10 days. Normal data is observed during the first 5 days, while anomalies are injected into the remaining 5 days. Secure Water Treatment (SWaT) [^3] https://itrust.sutd.edu.sg/testbeds/secure-water-treatment-swat/ (Mathur and Tippenhauer, 2016) contains data from a water treatment platform monitored by 51 sensors over 11 days. A total of 41 anomalies were deliberately generated during the last 4 days using various attack scenarios. SWAN [^4] https://dataverse.harvard.edu/api/access/datafile/:persistentId?persistentId=doi:10.7910/DVN/EBCFKM/K9AOSI (Angryk et al., 2020) is a publicly accessible, MTS benchmark derived from solar photospheric vector magnetograms in the Spaceweather HMI Active Region Patch series. GECCO [^5] http://www.spotseven.de/gecco/gecco-challenge/gecco-challenge-2018/ (Moritz et al., 2018) is a dataset on drinking water quality adapted for applications of the ‘Internet of Things’. It was introduced during the 2018 Genetic and Evolutionary Computation Conference.
In addition to the above MTS datasets, we also evaluated UTS datasets: UCR [^6] https://www.cs.ucr.edu/~eamonn/time_series_data_2018/ (Wu and Keogh, 2021) was provided for the Multi-dataset Time Series Anomaly Detection Competition at KDD2021. This dataset includes 250 UTS subdatasets, each containing one anomaly within its subsequences. Yahoo [^7] https://webscope.sandbox.yahoo.com/catalog.php?datatype=s&did=70 (Laptev et al., 2015) consists of 367 hourly time series with labels. Our focus is on the A1 benchmark, which includes 67 UTS representing “real” production traffic data from Yahoo properties. Key Performance Indicators (KPI) [^8] https://github.com/NetManAIOps/KPI-Anomaly-Detection are sourced from real-world Internet company scenarios. It contains service and machine KPIs, such as response time, page views, CPU usage, and memory usage.
For all datasets, we adopt their original train/test splits, with the training data being unlabeled, except for the Yahoo and GECCO datasets, where we adopt a 50/50 split for training and testing.
Appendix B List of Symbols and Notations
You can find the list of symbols and notations, and their definitions in Table 5.
Table 5. List of Symbols and Notations
Symbol Definition
$\mathcal{D}$ Time series dataset
$T$ Number of time steps in a time series window
$D$ Number of time series dimensions
$\mathbf{X}$ Time series window of shape $T\times D$
$\widetilde{\mathbf{X}}$ Generated anomalous time series sample
$\hat{\mathbf{X}}$ Reconstructed normal sample from $\mathbf{z}$
$\Gamma$ Generative mapping from normal to anomalous samples
$\phi$ Parameters of the encoder
$\theta$ Parameters of the decoder
$\mathbf{z}$ Latent representation of a normal sample
$\widetilde{\mathbf{z}}$ Perturbed latent representation of an anomaly
$L$ Latent space dimensionality
$p(\mathbf{X})$ Normal data distribution
$p(\widetilde{\mathbf{X}})$ Anomalous data distribution
$q(\mathbf{z}|\mathbf{X})$ Latent distribution of normal data
$p(\widetilde{\mathbf{z}})$ Shifted latent distribution for anomalies
$\mu$ Mean of the latent distribution
$\sigma$ Standard deviation of the latent distribution
$\sigma_{\text{prior}}$ Prior standard deviation for KL divergence loss
$\psi$ Learned perturbation scale
$\delta_{\text{min}}$ Minimum margin for anomaly distinctness
$\delta_{\text{max}}$ Maximum allowable deviation for realism
$\tau$ Threshold for deviation-based patching
Appendix C Algorithms Overview
The pseudo-code for GenIAS is provided in Algorithm 1, while the deviation-based patching approach is detailed in Algorithm 2.
Input: TS dataset $\mathcal{D}$, prior standard deviation $\sigma_{\text{prior}}$, loss coefficients $\alpha$, $\beta$, $\gamma$, $\zeta$.
Output: Trained encoder parameters $\phi$, decoder parameters $\theta$, perturbation scale $\psi$.
1Initialize encoder $q_{\phi}$, decoder $g_{\theta}$, latent space parameters $\mu$, $\sigma$, and perturbation scale $\psi$;
2 while training not converged do
      3 Sample a batch $\{\mathbf{\text{X}}_{i}\}_{i=1}^{B}\subset\mathcal{D}$ of size $B$;
      4 Encode: Obtain latent parameters $\mu_{i}$, $\sigma_{i}$ by $q_{\phi}(\mathbf{\text{X}}_{i})$;
      5 Sample latent vectors $\mathbf{\text{z}}_{i}\sim\mathcal{N}({\mu}_{i},\operatorname{diag}({\sigma}_{i}^{2}))$;
      6 Generate perturbed latent vectors $\widetilde{\mathbf{z}}_{i}$triangleright Eq. (1)
      7 Decode reconstructed samples: $\hat{\mathbf{X}}_{i}=g_{\theta}(\mathbf{z}_{i})$;
      8 Decode anomalous samples: $\hat{\mathbf{X}}_{i}\ =g_{\theta}(\widetilde{\mathbf{z}}_{i})$;
      9 Compute the reconstruction loss $\mathcal{L}_{\text{recon}}$; $\triangleright$ Eq. (2)
      10 Compute the perturbation loss $\mathcal{L}_{\text{perturb}}$; $\triangleright$ Eq. (3)
      11 Compute the zero-perturb loss $\mathcal{L}_{\text{zero-perturb}}$; $\triangleright$ Eq. (4)
      12 Compute the enhanced KL loss $\mathcal{L}_{\text{en-KL}}$; $\triangleright$ Eq. (5)
      13 $\mathcal{L}_{\text{total}}=\alpha\cdot\mathcal{L}_{\text{recon}}+\beta\cdot\mathcal{L}_{\text{perturb}}+\gamma\cdot\mathcal{L}_{\text{zero-perturb}}+\zeta\cdot\mathcal{L}_{\text{en-KL}}$;
      14 Update the model parameters $\phi$, $\theta$ and $\psi$  by minimizing $\mathcal{L}_{\text{total}}$;
15return $\phi$, $\theta$, $\psi$;
Algorithm 1 End-to-End Training of GenIAS
Input: Pretrained GenIAS model’s encoder $q_{\phi}$, decoder $g_{\theta}$ and perturbation scales $\psi$, time series window $\mathbf{X}$, threshold scaling factor $\tau$.
Output: Patched anomalous time series windows $\tilde{\mathbf{X}}_{\text{patched}}$.
Phase 1: Anomaly Generation
1 Load encoder $q_{\phi}$, decoder $g_{\theta}$, and perturbation scales $\psi$.
2Encode: Obtain latent parameters $\mu$, $\sigma$ by $q_{\phi}(\mathbf{\text{X}})$;
3Generate perturbed latent vectors $\widetilde{\mathbf{z}}$triangleright Eq. (1)
4Decode $\widetilde{\mathbf{z}}$ to generate preliminary anomalies: $\tilde{\mathbf{X}}=g_{\theta}(\tilde{\mathbf{z}})$;
Phase 2: Deviation-Based Patching
5 foreach dimension $d$ in $\mathbf{X}$ do
      6 $\text{deviation}_{d}=\left|\mathbf{X}_{d}-\tilde{\mathbf{X}}_{d}\right|^{2}$;
      7 $\text{amplitude}_{d}=\max(\mathbf{X}_{d})-\min(\mathbf{X}_{d})$;
      8 Generate patched output for dimension $d$:
$\tilde{\mathbf{X}}_{\text{patched},d}=\begin{cases}\tilde{\mathbf{X}}_{d}&\text{if deviation}_{d}>\tau\cdot\text{amplitude}_{d},\\ \mathbf{X}_{d}&\text{otherwise}.\end{cases}$
9return $\tilde{\mathbf{X}}_{\text{patched}}$ as the patched anomalous TS windows.
Algorithm 2 Patching with GenIAS
Appendix D Proofs of Theorem
D.1. Notation and Setup
Assume that the latent encoder produces a Gaussian posterior
(7) $q_{\phi}(z\mid{\mathbf{X}})=\mathcal{N}\left(\mu,\operatorname{diag}(\sigma^{2})\right),$
and that the prior is chosen as
(8) $p(z)=\mathcal{N}\left(0,\sigma^{2}_{\text{prior}}I\right),$
where $\sigma^{2}_{\text{prior}}$ is a hyperparameter and $\mu$ is a vector $\mu=(\mu_{1},...,\mu_{L})$. In the following, we consider the effect of using a prior $\sigma^{2}_{\text{prior}}<1$ in our Enhanced KL Loss (Equation 5).
D.2. Prerequisites
Lemma 0 (Compact Latent Space in Enhanced KL Loss).
Consider the KL divergence between the estimated posterior and the prior distributions for a latent dimension $j$. If the enhanced KL loss -Equation 5- is minimized (or its gradient dominates locally) with a prior variance $\sigma^{2}_{\text{prior}}<1$, then for each latent variable $z_{j}$ the optimal posterior variance satisfies $\sigma^{2}_{j}=\sigma^{2}_{\text{prior}}$. Hence, the variance of latent representations for normal samples are forced to be less than 1, i.e., more “compact”.
Proof.
For a one-dimensional Gaussian posterior $\mathcal{N}(\mu,\sigma^{2})$ and a prior $\mathcal{N}(0,\sigma^{2}_{\text{prior}})$, the KL divergence is
(9) $D_{\text{KL}}(q(z\mid\mathbf{X})\parallel p(z))=\frac{1}{2}\left(\frac{\sigma^{2}}{\sigma^{2}_{\text{prior}}}+\frac{\mu^{2}}{\sigma^{2}_{\text{prior}}}-1-\log\frac{\sigma^{2}}{\sigma^{2}_{\text{prior}}}\right).$
If we minimize the KL term with respect to $\sigma^{2}$ (ignoring the contribution from the reconstruction term and assuming $\mu$ is fixed), we define a function $f$ that takes the part that depends on $\sigma^{2}$:
(10) $f(\sigma^{2})=\frac{1}{2}\left(\frac{\sigma^{2}}{\sigma^{2}_{\text{prior}}}-\log\frac{\sigma^{2}}{\sigma^{2}_{\text{prior}}}\right).$
Taking the derivative,
(11) $\frac{df}{d\sigma^{2}}=\frac{1}{2}\left(\frac{1}{\sigma^{2}_{\text{prior}}}-\frac{1}{\sigma^{2}}\right).$
Setting the derivative to zero yields
(12) $\frac{1}{\sigma^{2}_{\text{prior}}}-\frac{1}{\sigma^{2}}=0\quad\Longrightarrow\quad\sigma^{2}=\sigma^{2}_{\text{prior}}.$
Thus, when $\sigma^{2}_{\text{prior}}<1$, the optimal posterior variance (in the sense of minimizing the KL divergence) is less than 1, leading to a more compact latent representation. ∎
Lemma 0 (Reduced Reconstruction Error for Normal Samples).
Assume that the decoder $g_{\theta}(z)$ is trained on normal data. If the latent representations of normal samples are forced to be compact (i.e. the variance $\sigma^{2}(\mathbf{X})$ is reduced), then a first-order Taylor expansion argument shows that the additional error due to sampling variability is reduced, thereby lowering the expected reconstruction error.
Proof.
For an input $\mathbf{X}$ with latent variable $z\sim q_{\phi}(z\mid\mathbf{X})$ and decoder $g_{\theta}(z)$, the reconstruction error is given by
(13) $\text{MSE}(\mathbf{X})=\mathbb{E}_{z\sim q_{\phi}(z\mid\mathbf{X})}\left[\|\mathbf{X}-g_{\theta}(z)\|^{2}\right].$
A first-order Taylor expansion of $g_{\theta}(z)$ around the mean $\mu(\mathbf{X})$ gives
(14) $g_{\theta}(z)\approx g_{\theta}(\mu(\mathbf{X}))+J_{g_{\theta}}(\mu(\mathbf{X}))(z-\mu(\mathbf{X})),$
where $J_{g_{\theta}}$ is the Jacobian of $g_{\theta}$ at $\mu(\mathbf{X})$. We substitute Equation 14 in Equation 13. While $\mu(\mathbf{X})$ stays the same in both Enhanced KL and KL cases, $z-\mu(\mathbf{X})$ varies in these two cases. Hence we focus on the variable term in this substitution, i.e., $\mathbb{E}\left[\|J_{g_{\theta}}(z-\mu(\mathbf{X}))\|^{2}\right]$. In a multi-dimensional latent space, if the posterior covariance is $\Sigma(\mathbf{X})=\operatorname{diag}(\sigma_{1}^{2}(\mathbf{X}),\dots,\sigma_{d}^{2}(\mathbf{X}))$, then the second-order contribution to the reconstruction error is approximately
(15) $\mathbb{E}\left[\|J_{g_{\theta}}(z-\mu(\mathbf{X}))\|^{2}\right]=\operatorname{trace}\left(J_{g_{\theta}}^{T}J_{g_{\theta}}\Sigma(\mathbf{X})\right).$
When the enhanced KL loss forces each $\sigma_{j}^{2}(\mathbf{X})$ to be small (namely $\sigma^{2}_{\text{prior}}<1$), the contribution of the variability term
(16) $\operatorname{trace}\left(J_{g_{\theta}}^{T}J_{g_{\theta}}\Sigma(\mathbf{X})\right)$
is reduced. Hence, the overall reconstruction error decreases. Figure 8 shows the impact of $\sigma^{2}_{\text{prior}}$ on the reconstruction MSE distribution for entity C-1 in MSL dataset. ∎
D.3. Proof of Theorem 1
Statement: Let a VAE be trained on normal data with a latent prior $\mathcal{N}(0,\sigma^{2}_{\text{prior}})$, where $\sigma_{\text{prior}}<1$. The encoder posterior for normal samples follows $\mathcal{N}(\mu,\sigma^{2}_{\text{normal}})$, with compactness enforced by KL regularization. If the encoder $\phi$ applies a perturbation scale $\psi$, inflating variance as $\sigma^{2}_{\text{anom}}=\psi\sigma^{2}_{\text{normal}},\quad\text{with }\psi>1,$ then the KL divergence between normal and anomalous latent distributions strictly increases compared to the case where $\sigma_{\text{prior}}\geq 1$.
Proof.
For generated anomalous samples, assume that the latent posterior is
(17) $q_{\text{normal}}(z)=\mathcal{N}(\mu_{\text{normal}},\operatorname{diag}(\sigma^{2}_{\text{normal}})),$
with $\sigma^{2}_{\text{normal}}\approx\sigma^{2}_{\text{prior}}$ (by Lemma 1). For anomalous inputs, suppose the encoder yields an inflated variance so that
(18) $q_{\text{anom}}(z)=\mathcal{N}(\mu_{\text{anom}},\operatorname{diag}(\psi\sigma^{2}_{\text{normal}})),$
with $\psi>1$. We assume that the means are similar, i.e. $\mu_{\text{anom}}\approx\mu_{\text{normal}}$, so that the primary difference is in the variances.
We aim to show that the KL divergence between the latent distributions of normal and anomalous samples strictly increases when the VAE prior variance is compact ($\sigma_{\text{prior}}<1$) compared to when no prior compactness is enforced ($\sigma_{\text{prior}}\geq 1$).
From the given equation the KL divergence between the anomalous and normal posterior distributions in one dimension is:
(19) $\begin{split}D_{\text{KL}}\left(\mathcal{N}(\mu,\psi\sigma^{2}_{\text{normal}})\Big{\|}\mathcal{N}(\mu,\sigma^{2}_{\text{prior}})\right)=\\ \frac{1}{2}\left(\frac{\psi\sigma^{2}_{\text{normal}}}{\sigma^{2}_{\text{prior}}}-1-\log\frac{\psi\sigma^{2}_{\text{normal}}}{\sigma^{2}_{\text{prior}}}\right).\end{split}$
We compare two scenarios:
  • Scenario A (Compact Latent Space): $\sigma_{\text{prior}}<1$, let $\sigma_{\text{prior}}=\alpha$, where $10<\alpha<$ 1.
  • Scenario B (Non-Compact Latent Space): $\sigma_{\text{prior}}\geq 1$, specifically, we consider $\sigma_{\text{prior}}=1$.
For Scenario A ($\sigma_{\text{prior}}=\alpha$):
(20) $D_{\text{KL},A}=\frac{1}{2}\left(\frac{\psi\sigma^{2}_{\text{normal}}}{\alpha^{2}}-1-\log\frac{\psi\sigma^{2}_{\text{normal}}}{\alpha^{2}}\right).$
For Scenario B ($\sigma_{\text{prior}}=1$):
(21) $D_{\text{KL},B}=\frac{1}{2}\left(\frac{\psi\sigma^{2}_{\text{normal}}}{1}-1-\log\frac{\psi\sigma^{2}_{\text{normal}}}{1}\right).$
Since $10<\alpha<$ 1, we observe that the fraction $\frac{\psi\sigma^{2}_{\text{normal}}}{\alpha^{2}}$ in $D_{\text{KL},A}$ is strictly larger than $\frac{\psi\sigma^{2}_{\text{normal}}}{1}$ in $D_{\text{KL},B}$.
The function $f(x)=x-1-\log x$ is monotonically increasing for $x>1$, meaning:
(22) $f\left(\frac{\psi\sigma^{2}_{\text{normal}}}{\alpha^{2}}\right)>f\left(\frac{\psi\sigma^{2}_{\text{normal}}}{1}\right).$
Thus, we obtain:
(23) $D_{\text{KL},A}>D_{\text{KL},B}.$
Since $D_{\text{KL},A}>D_{\text{KL},B}$, we conclude that the KL divergence between normal and anomalous latent distributions is strictly larger when the VAE is trained with a compact latent space ($\sigma_{\text{prior}}<1$).
This means that a compact latent space amplifies the separation between normal and anomalous samples under perturbation.
Appendix E Details of Evaluation Metrics
To support a consistent and meaningful evaluation of generated anomalies, we utilize a Deep SVDD (Ruff et al., 2018) representation with a fixed 128-dimensional embedding for all models. This representation provides a structured feature space. Additionally, we evaluate models using test datasets that contain labeled anomalies, allowing for a more reliable assessment of anomaly generation quality. Let $\Omega$ denote the DeepSVDD model. The representation of real anomalous samples in the test dataset is given by $V_{\text{real}}=\Omega(\text{X}_{\text{real anomalies}})$, while the representation of generated anomalous samples, produced by the anomaly generator model from the train dataset, is given by $V_{\text{gen}}=\Omega(\tilde{{\text{X}}})$.
E.1. Anomalous Representation Proximity (ARP)
ARP measures how well the generated anomalies cover the distribution of real anomalies in the representation space. For each real anomaly, we compute the distance to its closest generated anomaly. The inverse of the average of these minimum distances reflects how well the generated anomalies span the real anomaly space. Formally, if $V_{\text{real}}$ and $V_{\text{gen}}$ are the sets of real and generated anomaly representations, the ARP is defined as:
(24) $\text{ARP}=\frac{1}{1+\frac{1}{|V_{\text{real}}|}\sum_{\mathbf{v}^{r}\in V_{\text{real}}}\min_{\mathbf{v}^{g}\in V_{\text{gen}}}d(\mathbf{v}^{r},\mathbf{v}^{g})}.$
where $d(\cdot,\cdot)$ is the Euclidean distance metric. A higher ARP indicates that the generated anomalies are more similar to real anomalies in the representation space, meaning better realism.
E.2. Entropy-Based Diversity Index (EDI)
EDI quantifies how well a model generates a diverse set of anomalies by evaluating their distribution across the representation space. First, we collect the representations of all generated anomalies from all models into a global set:
(25) $V_{\text{all}}=\bigcup_{m=1}^{M}V_{\text{gen}}^{(m)},$
where $V_{\text{gen}}^{(m)}=\{\mathbf{v}_{1}^{(m)},\mathbf{v}_{2}^{(m)},\dots,\mathbf{v}_{N_{m}}^{(m)}\}$ represents the set of generated anomalies for model $m$, and $M$ is the number of models being evaluated. Next, we partition the overall representation space into $K$ non-overlapping regions $\{R_{1},R_{2},\dots,R_{K}\}$ based on $V_{\text{all}}$, ensuring that each anomaly belongs to exactly one region.
For a specific model $m$, the proportion of its generated anomalies in region $R_{i}$ is defined as:
(26) $p_{i}^{(m)}=\frac{|\{\mathbf{v}_{j}^{(m)}\in V_{\text{gen}}^{(m)}\mid\mathbf{v}_{j}^{(m)}\in R_{i}\}|}{N_{m}},$
where $N_{m}=|V_{\text{gen}}^{(m)}|$ is the total number of generated anomalies for model $m$, and ||$\cdot$|| ⋅ | denotes the cardinality of a set.
The EDI for model $m$ is then computed using Shannon Entropy (Shannon, 1948) as:
(27) $\text{EDI}^{(m)}=-\sum_{i=1}^{K}p_{i}^{(m)}\log p_{i}^{(m)}.$
A higher EDI indicates that a model generates anomalies that are well-distributed across the space, reflecting greater diversity rather than being concentrated in a few regions.
E.3. TSAD Evaluation Metrics
We use Best F1 score to measures the performance of TSAD models at the optimal threshold. AUPR and AUROC are holistic measures that assess model performance across various thresholds. All these metrics account for class imbalance, with AUPR placing greater emphasis on anomaly class performance, making it more reflective of anomaly detection capabilities.
In addition, the two recent metrics are chosen to evaluate TSAD performance at finer granularities: Affiliation F1 (derived from Affiliation-Precision and Affiliation-Recall) and PATE (Proximity-Aware Time series anomaly Evaluation that incorporates proximity-based weighting and buffer zones). Both metrics assess early or delayed detections.
Appendix F Implementation Details
All evaluations were performed on a system with an A40 GPU, 13 CPU cores, and 250 GB of RAM. We use the same hyperparameters for GenIAS across UTS and MTS datasets, as detailed in Table 6.
Table 6. hyperparameter configurations used in GenIAS.
Parameter Value
$T$ 200
$L$ UTS = 50, MTS = 100
$\sigma_{\text{prior}}$ 0.5
lr initial=$10^{-4}$, with scheduled reduction
epochs max=1000
kernel size 3
dropout 0.1
batch size 100
$\alpha$ 1.0
$\beta$ 0.1
$\gamma$ UTS = 0.0, MTS = 0.01
$\zeta$ 0.1
Appendix G Additional Results
G.1. Visualizing Enhanced KL Loss Effects
Separation of reconstruction error in the input space is illustrated in Figure 8.
Figure 8. Separation of reconstruction error (anomaly score) in the input space with a prior of 0.5 (left) and without a prior (right) on entity C-1 from the MSL dataset. The y-axis is logarithmic.
G.2. TSAD additional results
G.2.1. Critical Difference diagrams
Statistical significance in critical difference digrams was determined using a Friedman test followed by a post-hoc Wilcoxon signed-rank test with Holm correction (Demšar, 2006). The critical difference diagrams in Figure 9 show that GenIAS consistently ranks first across both multivariate (MTS) and univariate (UTS) time series settings for AUPR and F1 Score. In the MTS setting, GenIAS demonstrates a statistically significant improvement over other baselines, indicating its superior performance in capturing anomalies.
Figure 9. Critical difference diagrams for F1 and AUPR scores: (a, b) represent multivariate models, while (c, d) showcase univariate models.
G.2.2. TSAD performance in the other metrics.
In Table 7, we present the TSAD performance of GenIAS and the baseline methods in terms of three additional metrics: Affiliation F1 (Aff), PATE, and AUROC. In general, similar observations can be drawn as in Section 4.7 - GenIAS achieves superior performance compared to the baseline methods. Specifically, for Aff and PATE, GenIAS achieves the best overall performance across both MTS and UTS benchmark datasets. In terms of AUROC, GenIAS achieves the second-best overall performance and best overall performance among neural network-based methods, while Isolation Forest, a well-known legacy shallow AD method, yields surprisingly good performance.
Table 7. TSAD performance of GenIAS and baseline methods in terms of Aff, PATE and AUROC, with the best model in bold and the second-best model underlined. U and M denote results unavailable for models limited to UTS or MTS, respectively.
Multivariate Univariate Metric Cat. Methods MSL SMAP SMD SWAT GECCO SWAN Avg-M Yahoo KPI UCR Avg-U Avg Aff Classic OCSVM (1999) 0.742 0.761 0.733 0.197 0.425 0.711 0.595 0.923 0.761 0.710 0.798 0.663 LOF (2000) 0.775 0.762 0.666 0.613 0.227 0.745 0.631 0.906 0.674 0.718 0.766 0.676 IForest (2008) 0.719 0.728 0.699 0.535 0.259 0.824 0.627 0.907 0.760 0.686 0.784 0.680 PU LSTM-VAE (2018) 0.824 0.782 0.630 0.324 0.312 0.726 0.600 M M M M M US-WI Donut (2018) U U U U U U U 0.777 0.682 0.685 0.715 - Omni (2019) 0.680 0.779 0.706 0.156 0.230 0.608 0.527 M M M M M THOC (2020) 0.778 0.781 0.677 0.679 0.321 0.713 0.658 0.796 0.676 0.642 0.705 0.674 MTAD-GAT(2020) 0.833 0.848 0.731 0.355 0.308 0.745 0.637 M M M M M AnomalyTran (2021) 0.774 0.766 0.664 0.482 0.551 0.537 0.629 M M M M M TranAD (2022) 0.771 0.751 0.719 0.606 0.233 0.661 0.624 0.917 0.779 0.698 0.798 0.682 TS2Vec (2022) 0.694 0.695 0.750 0.700 0.520 0.728 0.681 0.867 0.818 0.671 0.785 0.716 TimesNet (2023) 0.769 0.710 0.795 0.693 0.829 0.710 0.751 0.917 0.844 0.712 0.824 0.775 Dcdetector (2023) 0.706 0.680 0.528 0.693 0.671 0.710 0.665 0.737 0.666 0.671 0.691 0.674 US-I NCAD (2022) 0.685 0.664 0.732 0.693 0.887 0.710 0.729 0.645 0.753 0.747 0.715 0.724 CutAddPaste (2024) 0.762 0.753 0.693 0.764 0.659 0.875 0.751 0.832 0.818 0.820 0.823 0.775 COUTA (2024) 0.785 0.762 0.764 0.645 0.592 0.731 0.713 0.906 0.851 0.730 0.829 0.752 CARLA (2025) 0.816 0.808 0.733 0.592 0.756 0.932 0.773 0.874 0.701 0.805 0.793 0.780 Ours GenIAS 0.851 0.856 0.731 0.596 0.763 0.935 0.789 0.921 0.788 0.814 0.841 0.806 PATE Classic OCSVM (1999) 0.267 0.270 0.363 0.698 0.191 0.704 0.416 0.780 0.343 0.033 0.385 0.405 LOF (2000) 0.242 0.251 0.319 0.187 0.042 0.706 0.291 0.690 0.159 0.056 0.302 0.295 IForest (2008) 0.195 0.197 0.306 0.743 0.191 0.799 0.405 0.735 0.325 0.027 0.362 0.391 PU LSTM-VAE (2018) 0.318 0.278 0.436 0.693 0.061 0.753 0.423 M M M M M US-WI Donut (2018) U U U U U U U 0.358 0.120 0.018 0.165 - Omni (2019) 0.143 0.162 0.422 0.702 0.076 0.709 0.369 M M M M M THOC (2020) 0.307 0.218 0.147 0.508 0.104 0.563 0.308 0.488 0.317 0.016 0.274 0.296 MTAD-GAT(2020) 0.391 0.371 0.445 0.115 0.247 0.707 0.379 M M M M M AnomalyTran (2021) 0.267 0.283 0.332 0.683 0.112 0.685 0.394 M M M M M TranAD (2022) 0.238 0.279 0.458 0.192 0.141 0.688 0.333 0.777 0.408 0.025 0.403 0.356 TS2Vec (2022) 0.143 0.150 0.137 0.147 0.055 0.530 0.194 0.636 0.247 0.144 0.342 0.243 TimesNet (2023) 0.282 0.207 0.452 0.083 0.240 0.636 0.317 0.760 0.355 0.042 0.386 0.340 Dcdetector (2023) 0.176 0.135 0.097 0.124 0.025 0.455 0.169 0.184 0.052 0.025 0.087 0.141 US-I NCAD (2022) 0.157 0.161 0.129 0.107 0.272 0.556 0.230 0.145 0.173 0.130 0.149 0.203 CutAddPaste (2024) 0.269 0.347 0.200 0.613 0.252 0.842 0.421 0.237 0.741 0.262 0.413 0.418 COUTA (2024) 0.270 0.244 0.447 0.159 0.285 0.756 0.360 0.684 0.335 0.054 0.358 0.359 CARLA (2025) 0.493 0.413 0.488 0.655 0.193 0.911 0.526 0.695 0.378 0.301 0.458 0.503 Ours GenIAS 0.571 0.516 0.553 0.718 0.274 0.941 0.596 0.838 0.456 0.331 0.542 0.578 AUC-ROC Classic OCSVM (1999) 0.605 0.634 0.762 0.799 0.446 0.714 0.660 0.918 0.728 0.567 0.738 0.686 LOF (2000) 0.610 0.584 0.708 0.660 0.370 0.789 0.620 0.890 0.604 0.575 0.690 0.643 IForest (2008) 0.571 0.579 0.768 0.842 0.927 0.862 0.758 0.915 0.749 0.570 0.745 0.754 PU LSTM-VAE (2018) 0.618 0.650 0.771 0.806 0.506 0.763 0.686 M M M M M US-WI Donut (2018) U U U U U U U 0.676 0.640 0.498 0.605 - Omni (2019) 0.387 0.570 0.794 0.811 0.430 0.696 0.615 M M M M M THOC (2020) 0.654 0.604 0.678 0.318 0.395 0.578 0.538 0.873 0.799 0.558 0.743 0.606 MTAD-GAT(2020) 0.674 0.641 0.812 0.822 0.590 0.749 0.715 M M M M M AnomalyTran (2021) 0.641 0.653 0.764 0.809 0.698 0.678 0.707 M M M M M TranAD (2022) 0.632 0.645 0.808 0.677 0.519 0.677 0.660 0.889 0.763 0.515 0.722 0.681 TS2Vec (2022) 0.503 0.517 0.496 0.548 0.501 0.573 0.523 0.698 0.567 0.501 0.589 0.545 TimesNet (2023) 0.664 0.578 0.837 0.243 0.924 0.598 0.641 0.892 0.750 0.597 0.746 0.676 Dcdetector (2023) 0.535 0.519 0.579 0.504 0.504 0.499 0.523 0.603 0.504 0.509 0.539 0.528 US-I NCAD (2022) 0.491 0.493 0.572 0.365 0.692 0.592 0.534 0.567 0.671 0.609 0.616 0.561 CutAddPaste (2024) 0.673 0.631 0.631 0.777 0.789 0.464 0.661 0.684 0.810 0.617 0.704 0.675 COUTA (2024) 0.661 0.606 0.811 0.643 0.660 0.802 0.697 0.911 0.787 0.559 0.752 0.716 CARLA (2025) 0.694 0.677 0.657 0.783 0.696 0.510 0.670 0.543 0.604 0.735 0.627 0.655 Ours GenIAS 0.745 0.718 0.633 0.786 0.811 0.536 0.705 0.778 0.691 0.778 0.749 0.720
(a) KPI - Entity c69a50cf-ee03-3bd7-831e-407d36c7ee91
(b) SMD - Entity machine-2-3
(c) UCR - Entity 222_UCR_Anomaly_mit14046longtermecg_56123_91200_91700
Figure 10. Each plot visualises a different time series (TS) injection or generation method using t-SNE. Blue markers represent the original anomalous TS, while orange markers represent the injected or generated anomalous TS.
G.2.3. Additional Visualizations
To further illustrate the realistic anomaly generation of GenIAS, we present three additional visualizations comparing generated and real anomalies from both GenIAS and the US-i baselines in Figure 10. The observations are consistent with Section 4.6 in the main text, demonstrating that GenIAS produces comparatively more realistic anomalies with better smoothness in the latent space.

### v2

@@
 GenIAS: Generator for Instantiating Anomalies in Time Series
-Abstract.
-A recent and promising approach for building time series anomaly detection (TSAD) models is to inject synthetic samples of anomalies within real data sets.
-The existing injection mechanisms have significant limitations — most of them rely on ad hoc, hand-crafted strategies which fail to capture the natural diversity of anomalous patterns, or are restricted to univariate time series settings.
-To address these challenges, we design a generative model for TSAD using a variational autoencoder, which is referred to as a Generator for Instantiating Anomalies in Time Series (GenIAS).
-GenIAS is designed to produce diverse and realistic synthetic anomalies for TSAD tasks.
-By employing a novel learned perturbation mechanism in the latent space and injecting the perturbed patterns in different segments of time series, GenIAS can generate anomalies with greater diversity and varying scales.
-Further, guided by a new triplet loss function, which uses a min-max margin and a new variance-scaling approach to further enforce the learning of compact normal patterns, GenIAS ensures that anomalies are distinct from normal samples while remaining realistic.
-The approach is effective for both univariate and multivariate time series.
-We demonstrate the diversity and realism of the generated anomalies.
-Our extensive experiments demonstrate that GenIAS — when integrated into a TSAD task — consistently outperforms seventeen traditional and deep anomaly detection models, thereby highlighting the potential of generative models for time series anomaly generation.
-Anomaly generator, Time series, Anomaly detection
-1. Introduction
-Unsupervised methods for Time Series Anomaly Detection (TSAD) often learn a model for boundary of normal data without access to labeled anomalies.
-This reliance on normal data makes models vulnerable to unseen or diverse anomaly patterns, which are common in time series settings.
-This is particularly true in complex multivariate scenarios, where anomalous patterns may vary significantly across different dimensions or even be correlated.
-Anomaly injection techniques are an increasingly popular approach to address the limitations of solely relying on normal-boundary learning.
-Models like CARLA (Darban et al., 2025), CutAddPaste (Wang et al., 2024), NCAD (Carmona et al., 2022), and COUTA (Xu et al., 2024) have demonstrated the potential of synthetic anomaly injection in improving detection capabilities.
-However, these models face notable limitations.
-CARLA, CutAddPaste, and COUTA are restricted to predefined anomaly types, such as point, seasonal, trend, and shapelet anomalies.
-CutAddPaste further employs a rigid augmentation strategy (of “cut-add-paste”), which may not align well with realistic patterns in complex applications.
-NCAD, on the other hand, relies on prior knowledge about anomalies to generate realistic anomalies.
-However, it is limited by the anomalies present in the datasets and hence may not be able to generate diverse types of anomalies.
-Finally, all the existing anomaly injection models are susceptible to slight variations in the time series, as they generate anomalies in the raw time series space.
-We address these gaps by introducing the Generator for Instantiating Anomalies in Time Series (GenIAS), a novel generative model designed to produce diverse and realistic synthetic anomalies for TSAD.
-In contrast to previous approaches, GenIAS operates in the latent space rather than the raw time series space.
-This makes it less susceptible to minor variations in the raw time series and allows it to learn more generalized normal patterns.
-While generative models mainly used to generate normal data, we are able to leverage the power of variational neural models to generate anomalies that align well with the temporal properties of time series, and with correlations between variables if the input time series is multivariate.
-GenIAS builds upon a hybrid Temporal Convolutional Network-Variational Autoencoder (TCN-VAE) architecture to move beyond rigid, handcrafted augmentation strategies and maps the complex properties of time series data into a multidimensional latent space.
-By employing a learned perturbation mechanism applied to the latent space variance and injecting the perturbed patterns into different segments of the time series using a patching mechanism, GenIAS generates anomalies with varying scales and enhanced diversity, resulting in more realistic anomaly patterns.
-We use the term realism to refer to realistic anomaly patterns in this paper.
-Additionally, we employ an adjusted triplet loss function, guided by a min-max margin to guarantee that the generated anomalies are both distinct from normal data and realistic at the same time.
-Further, we introduce a novel variance-scaling approach, which enforces the compactness of latent representations, making the generated anomalies more distinguishable from normal samples.
-We demonstrate the effectiveness of GenIAS when extended as a TSAD model.
+Abstract
+Synthetic anomaly injection is a recent and promising approach for time series anomaly detection (TSAD), but existing methods rely on ad hoc, hand-crafted strategies applied to raw time series that fail to capture diverse and complex anomalous patterns, particularly in multivariate settings.
+We propose a synthetic anomaly generation method named Generator for Instantiating Anomalies in Time Series (GenIAS), which generates realistic and diverse anomalies via a novel learnable perturbation in the latent space of a variational autoencoder.
+This enables abnormal patterns to be injected across different temporal segments at varying scales based on variational reparameterization.
+To generate anomalies that align with normal patterns while remaining distinguishable, we introduce a learning strategy that jointly learns the perturbation scale and compact latent representations via a tunable prior, which improves the distinguishability of generated anomalies, as supported by our theoretical analysis.
+Extensive experiments show that GenIAS produces more diverse and realistic anomalies, and that detection models trained with these anomalies outperform 17 baseline methods on 9 popular TSAD benchmarks.
+Machine Learning, ICML
+1 Introduction
+Unsupervised TSAD methods often learn the boundary of normal data for detection without labeled anomalies, making them vulnerable to diverse or subtle anomaly patterns.
+This is particularly true in complex multivariate scenarios, where anomalous patterns can vary significantly across dimensions or involve intricate correlations.
+Anomaly injection techniques are gaining popularity for enhancing normal-boundary learning by addressing the lack of labels.
+Models like CARLA (Darban et al., 2025), CutAddPaste (Wang et al., 2024), NCAD (Carmona et al., 2022), and COUTA (Xu et al., 2024) show the promise of synthetic anomaly injection in improving detection, but each faces notable limitations.
+For instance, CARLA, CutAddPaste, and COUTA are limited to predefined anomaly types, such as point, seasonal, and shapelet anomalies.
+CutAddPaste employs a fixed “cut-add-paste” augmentation strategy, which may produce unrealistic patterns of time series.
+NCAD relies on prior knowledge of anomalies to generate realistic anomalies, but the diversity of its generated anomalies is constrained by the types of anomalies present in the training data.
+Moreover, most existing anomaly injection models operate directly in raw time series spaces, without modelling their normal patterns.
+This can result in unrealistic alignment between injected anomalies and the original time series components, producing anomalies that are out of context or entangled with normal patterns, and thus failing to provide anomaly discriminative information for TSAD training.
+We address these gaps by introducing GenIAS, a novel generative model designed to produce realistic and diverse synthetic anomalies for time series data.
+In contrast to existing methods that operate directly on raw time series, GenIAS performs anomaly generation in the latent space using a hybrid Temporal Convolutional Network–Variational Autoencoder (TCN-VAE) architecture.
+Through variational reparameterization, this design encodes the temporal and inter-variable dependencies of normal patterns in multivariate time series within a continuous latent space, enabling fine-grained and controllable perturbations for generating diverse and realistic anomalies.
+This is achieved by our novel learnable perturbation mechanism in the learned latent space, which modulates latent variance to induce anomalies with varying degrees of deviation across different time series components.
+To produce anomalies that align with the temporal properties of the original time series while remaining distinguishable from normal patterns, we introduce a joint training strategy that learns both the latent distribution of normal time series and the perturbation mechanism by enforcing two constraints.
+First, we propose a perturbation loss that encourages deviations of reasonable magnitude, avoiding perturbations that are either too trivial or excessively large.
+Second, we introduce a compact loss with a tunable variance prior initialized to a small value (e.g., <1<1) to induce compact latent representations in VAEs, as illustrated in Fig.
+1.
+Our theoretical analysis shows that such compactness reduces overlap between the generated anomalies and normal patterns when applying the proposed learnable perturbation.
+GenIAS injects anomalies through two stages, where a learned perturbation is first applied to the latent variance of normal time series representations to produce anomalous latent samples that are decoded into anomalous windows.
+To better reflect realistic TSAD scenarios, where anomalies are often localized rather than spanning all dimensions, a deviation-based patching strategy is then applied to selectively inject anomalous components from the anomalous window into the original window on a per-dimension basis, while preserving normal context.
+To evaluate anomaly generation quality, we introduce two complementary representation-based metrics that assess realism, defined as proximity to real anomalies in representation space, and diversity of the generated anomalies.
+GenIAS achieves a better balance in anomaly generation quality, and detection models trained with its generated anomalies consistently outperform 17 baselines across nine univariate and multivariate datasets.
 Our key contributions are as follows:
-  • We propose GenIAS, which is a novel anomaly generator designed to seamlessly handle both univariate time series (UTS) and multivariate time series (MTS).
-By employing deep generative techniques, GenIAS produces anomalies that are both diverse and realistic.
-  • We propose a novel loss function for our VAE-based anomaly generator, enforcing compact latent representations via a tailored prior variance.
-This improves the separation between normal and anomalous data in latent space, enhancing anomaly generation (see Figure 1).
-A theoretical analysis supports its effectiveness.
-  • We demonstrate GenIAS’s ability to generate diverse and realistic anomalies through two newly introduced metrics.
-The evaluation is conducted on four state-of-the-art TSAD models leveraging anomaly injection, across nine widely used MTS and UTS datasets.
-  • We demonstrate GenIAS’s broad applicability by integrating it into a state-of-the-art TSAD model and comparing it with 17 baselines.
-Extensive experiments on nine diverse UTS and MTS datasets confirm its effectiveness, with performance gains directly linked to the quality of GenIAS’s anomaly generation.
-Figure 1.
-An illustration of our key idea: Variance-scaling enforces latent space compactness in the latent space, strengthening the separation between normal and generated anomalous samples by GenIAS, in latent space.
-The bottom figure, with $\sigma_{\text{prior}}$ = 0.5, visually shows this effect.
-2. Related Work
-The field of time series anomaly detection has seen remarkable progress, and include traditional statistical methods as well as state-of-the-art deep learning approaches (Schmidl et al., 2022; Audibert et al., 2022; Zamanzadeh Darban et al., 2024).
-Early efforts included models like One-Class SVM (OC-SVM) (Schölkopf et al., 1999), LOF (Breunig et al., 2000), and Isolation Forest (Fei Tony Liu and Zhou, 2008); however, these methods faced challenges in handling high-dimensional and sequential data.
-Deep learning models like Donut (Xu et al., 2018), LSTM-VAE (Park et al., 2018) and OmniAnomaly (Su et al., 2019) introduced temporal dependency modeling and probabilistic approaches in TSAD.
-Representation-based methods, have pushed TSAD forward by focusing on meaningful embeddings.
-TS2Vec (Yue et al., 2022) applies self-supervised learning to capture multi-level semantic representations, while DCdetector (Yang et al., 2023) employs dual attention for permutation-invariant representations.
-Recent models such as TimesNet (Wu et al., 2023) and TranAD (Tuli et al., 2022a) model complex dependencies by leveraging attention mechanisms.
-Methods like MTAD-GAT (Zhao et al., 2020) utilize graph attention networks to model multivariate dependencies and THOC (Shen et al., 2020a) incorporates hierarchical modeling to capture temporal patterns at multiple granularities.
+Figure 1: Effect of compactness.
+A smaller prior $(\sigma_{\text{prior}}=0.5$ vs $\sigma_{\text{prior}}=1.0)$ variance produces compact normal latent representations and improved separation from generated anomalies
+  • We propose a novel synthetic anomaly generation method, GenIAS, that operates in the latent space of variational autoencoders and generates diverse anomalies with varying degrees of deviation across different time series components, validating anomaly injection via controllable latent space perturbation as a promising strategy for improving unsupervised TSAD.
+  • We introduce an optimization strategy for GenIAS that jointly learns the latent distribution of time series data and the perturbation mechanism.
+This ensures that injected deviations are informative and distinguishable while remaining coherent with the normal time series.
+  • We propose two metrics for evaluating the diversity and realism for measuring anomaly generation, where GenIAS demonstrates better and more balanced results compared to the synthetic anomaly generation methods.
+We further show that GenIAS-enabled detection models demonstrate superior performance compared to 17 baseline methods, translating improved anomaly generation into stronger TSAD performance.
+2 Related Work
+The field of TSAD has advanced significantly, spanning traditional statistical methods to SOTA deep learning approaches (Schmidl et al., 2022; Audibert et al., 2022; Zamanzadeh Darban et al., 2024).
+Early efforts included models like One-Class SVM (OC-SVM) (Schölkopf et al., 1999), LOF (Breunig et al., 2000), and Isolation Forest (Fei Tony Liu and Zhou, 2008), which struggled with handling high-dimensional and sequential data.
+In contrast, deep learning models like Donut (Xu et al., 2018), LSTM-VAE (Park et al., 2018) and OmniAnomaly (Su et al., 2019) introduced temporal dependency modeling, advancing TSAD performance.
+Representation-based methods have pushed TSAD forward by focusing on meaningful embeddings.
+TS2Vec (Yue et al., 2022) applies self-supervised learning to capture multi-level semantic features, while DCdetector (Yang et al., 2023) employs dual attention for permutation-invariant representations.
+Recent models such as TimesNet (Wu et al., 2023) and TranAD (Tuli et al., 2022b) leverage attention mechanisms to capture complex dependencies.
+Methods like MTAD-GAT (Zhao et al., 2020) utilize graph attention networks to model multivariate dependencies, and THOC (Shen et al., 2020a) incorporates hierarchical modelling to capture temporal patterns across multiple scales.
 Anomaly Transformer (Xu et al., 2021) leverages anomaly-sensitive features through attention mechanisms.
-OE (Hendrycks et al., 2019) integrates external data during training to expose the model to out-of-distribution instances, thereby making it more powerful.
-NCAD(Carmona et al., 2022) adapts OE for time series by infusing mixed contextual and random point anomalies to better separate normal and anomalous patterns.
-Anomaly injection and perturbation methods have emerged as crucial strategies for enhancing TSAD models.
-Approaches like CutAddPaste (Wang et al., 2024) and COUTA (Xu et al., 2024) simulate diverse anomaly patterns during training.
-These techniques have also been integrated into advanced models like CARLA (Darban et al., 2025) to create negative pairs for contrastive learning, enhancing generalization to unseen anomaly types.
-3. GenIAS Method
-3.1. Problem Statement
-Let $\mathcal{D}$ represent a time series dataset consisting of $N$ time series windows, where each window $\mathbf{X}\in\mathbb{R}^{T\times D}$ is a time series window of length $T$ with $D$ dimension(s).
-Assume that samples $X$ in $\mathcal{D}$ is drawn from an unknown data distribution $p(\mathbf{X})$.
-In TSAD datasets, the training data is typically unlabeled, and it is assumed normal since anomalies are rare and have negligible impact.
-The objective is to use a generative mapping $\Gamma:\mathbf{X}\rightarrow\widetilde{\mathbf{X}}$ to transform a normal sample $\mathbf{X}$ into an anomalous sample $\widetilde{\mathbf{X}}\in\mathbb{R}^{T\times D}$, such that the anomalous distribution $p(\widetilde{\mathbf{X}})$ satisfies $p(\widetilde{\mathbf{X}})\neq p(\mathbf{X})$.
-The anomalies must preserve essential temporal properties in the normal data while exhibiting deviations that make them statistically separated.
-The first challenge is to define a perturbation mechanism on the latent representation $\mathbf{z}\in\mathbb{R}^{L}$ with $L$ dimensions, parameterized by $p(\mathbf{z}\mid\mathbf{X})$, to produce a shifted latent distribution $p(\widetilde{\mathbf{z}})$.
-This ensures that the anomalies are distinguishable, satisfying $p(\widetilde{\mathbf{z}})\neq p(\mathbf{z})$ while retaining plausibility by remaining within the data manifold.
-Furthermore, by perturbing the latent representations, our model is less susceptible to the slight deviations (e.g., noise) in the raw time series space.
-The second challenge is to densify the normal latent distribution $p(\mathbf{z})$, enabling the mapping $\Gamma(\mathbf{X})=\hat{\mathbf{X}}$ to accurately reconstruct normal samples and establish a clear boundary for generating anomalies.
-Together, addressing these challenges guarantees that the generated anomalies achieve a balance between diversity and realism.
-This approach is designed to handle univariate ($D=1$) and multivariate ($D>1$) time series.
-3.2. Anomaly Generation
-GenIAS is a generative model designed to learn compact representations of normal time series data and generate realistic anomalies by strategically perturbing the latent space.
-The model comprises three main components: an encoder, a latent space representation, and a decoder.
-It is trained end-to-end with an objective function that balances reconstruction accuracy, latent space regularization, and anomaly generation quality.
-3.2.1. Model Workflow and Perturbation Mechanism
-The input to the model is a time series window $\mathbf{X}$ of shape $T\times D$.
-The input $\mathbf{X}$ is normalized to ensure consistency across features, improving model stability during training.
-The model architecture combines a TCN and a VAE to combine the strengths of both approaches.
-The VAE learns a latent representation of normal data and perturbs it in the latent space to enhance robustness to noisy time series and improve the learning of normal patterns.
-The TCN captures temporal dependencies within the time series more effectively.
-The encoder maps the input sample $\mathbf{X}$ into a latent representation $\mathbf{z}$ by capturing temporal dependencies through stacked convolutional layers.
-These layers, interleaved with activation functions and dropout, effectively model complex temporal patterns while reducing overfitting.
-The encoder outputs the parameters $\mu$ (mean) and $\sigma^{2}$ (variance), which define a Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$ in the latent space, where $\mu$ and $\sigma$ are vectors of size $L$, the same as the size of latent vector $\mathbf{z}$.
-The latent space representation forms the core of the model.
-For normal samples, latent vectors $\mathbf{z}$ are sampled as: $\mathbf{z}\sim\mathcal{N}(\mu,\sigma^{2})$.
-To generate anomalous representations $\tilde{\mathbf{z}}$, the standard deviation $\sigma$ is perturbed while keeping $\mu$ unchanged:
-(1) $\tilde{\mathbf{z}}=\mu+\psi\cdot(\sigma\odot\epsilon),$
-where $\psi$ is the learned perturbation scale, and $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is a noise vector.
-Perturbing $\sigma$ instead of $\mu$ ensures that the anomalies deviate in a way that preserves the central tendency of the normal data, thereby maintaining realism.
-Altering $\mu$ could disrupt this central tendency, leading to unrealistic deviations that do not align with the underlying data manifold.
-The decoder reconstructs the input sample from $\mathbf{z}$ and $\tilde{\mathbf{z}}$.
-Using transpose convolutional layers upscales the latent representations back to the original input dimensions while preserving the temporal structure.
-The decoder outputs the following:
-  • $\hat{\mathbf{X}}$: The reconstructed normal window, derived from $\mathbf{z}$.
-  • $\tilde{\mathbf{X}}$: The generated anomalous window, derived from $\tilde{\mathbf{z}}$.
-Figure 2 illustrates the overall workflow of the model, including the TCN-VAE architecture and the perturbation mechanism.
-The training process optimizes a combination of losses to ensure accurate reconstruction, realistic anomaly generation, and a compact latent space learning (see Algorithm 1 for details).
-Figure 2.
-The overall architecture of GenIAS: The model comprises an encoder, a structured latent space, and a decoder.
-The encoder, built with a TCN-VAE, maps input time series windows into a latent representation by learning temporal dependencies.
-The latent space regularization ensures compact normal representations, while anomalies are generated by perturbing the variance of the learned distribution without altering the mean, preserving realistic deviations.
+OE (Hendrycks et al., 2019) improves robustness by introducing out-of-distribution examples during training.
+NCAD (Carmona et al., 2022) extends this idea to time series by injecting mixed contextual and random point anomalies to better separate normal and anomalous patterns.
+Anomaly injection and perturbation methods have become key strategies in TSAD, with methods like CutAddPaste (Wang et al., 2024) and COUTA (Xu et al., 2024) simulating diverse anomaly patterns during training.
+These techniques are also used in CARLA (Darban et al., 2025) to create negative pairs for contrastive learning, enhancing generalization to unseen anomalies.
+Figure 2: Overall architecture of GenIAS: The model comprises an encoder, a structured latent space, and a decoder.
+The encoder, based on a TCN-VAE, maps input time series windows into a latent representation by learning temporal dependencies.
+Latent space regularization encourages compact normal representations, and anomalies are generated by perturbing the variance of the learned distribution.
 The decoder reconstructs both normal and perturbed windows using transpose convolutional layers.
-3.2.2. Reconstruction Loss
-This portion of the overall loss is designed to accurately reconstruct normal input samples.
-It is defined using the Mean Squared Error (MSE) between an input sample $\mathbf{X}$ and its reconstruction $\hat{\mathbf{X}}$:
-(2) $\mathcal{L}_{\text{recon}}=\frac{1}{|\mathcal{D}|}\sum_{\mathbf{X}\in\mathcal{D}}\frac{1}{T\cdot D}\sum_{t=1}^{T}\sum_{d=1}^{D}\big{(}\mathbf{X}[t,d]-\hat{\mathbf{X}}[t,d]\big{)}^{2},$
-This loss penalizes the discrepancy between $\mathbf{X}$ and $\hat{\mathbf{X}}$, ensuring that the model captures the key temporal and structural features of normal samples.
-3.2.3. Perturbation Loss
-The perturbation loss ensures that the generated anomalies $\tilde{\mathbf{X}}$ are both distinct from normal samples $\mathbf{X}$ and exhibit realistic deviations.
-This is achieved through two complementary terms: a triplet loss to enforce distinctiveness and a regularization term to maintain realism.
-The first term, the triplet loss, encourages the original time series sample ${\mathbf{X}}$ to be closer to the reconstructed sample $\hat{\mathbf{X}}$ than the generated anomalous sample $\tilde{\mathbf{X}}$, ensuring that the anomalies are sufficiently distinct from original time series.
-By maximizing the margin between $d(\mathbf{X},\hat{\mathbf{X}})$ and $d(\mathbf{X},\tilde{\mathbf{X}})$, where $d(x,y)$ denotes the MSE distance, this term pushes the anomalous samples further away from the normal samples in the raw time series space.
-We also ensure that this margin is at least $\delta_{\text{min}}$, thus the generated anomalies be more distinct.
-The second term (the regularization term) prevents the anomalous samples $\tilde{\mathbf{X}}$ from deviating excessively from the original samples, thereby preserving realism.
-It penalizes deviations where distance $d(\mathbf{X},\tilde{\mathbf{X}})$ exceeds a maximum allowable threshold $\delta_{\text{max}}$.
-This ensures that the generated anomalies remain plausible and do not degenerate into unstructured or unrealistic noise.
-The perturbation loss is formulated as:
-(3) $\begin{split}\mathcal{L}_{\text{perturb}}=&\frac{1}{|\mathcal{D}|}\sum_{\mathbf{X}\in\mathcal{D}}\Big{[}\max\big{(}d(\mathbf{X},\hat{\mathbf{X}})-d(\mathbf{X},\tilde{\mathbf{X}})+\delta_{\text{min}},0\big{)}\Big{]}\\ &+\frac{1}{|\mathcal{D}|}\sum_{\mathbf{X}\in\mathcal{D}}\max\big{(}d(\mathbf{X},\tilde{\mathbf{X}})-\delta_{\text{max}},0\big{)},\end{split}$
-Together, these terms ensure that the generated anomalies achieve the dual objectives of being distinct yet realistic.
-3.2.4. Zero-Perturbation Loss
-The zero-perturbation loss focuses on dimensions in the input sample $\mathbf{X}$ that are entirely zero.
-It ensures that anomalies $\tilde{\mathbf{X}}$ introduce meaningful perturbations in these dimensions, encouraging deviations from the original sample where appropriate.
-The loss is defined as:
-(4) $\mathcal{L}_{\text{zero-perturb}}=\frac{1}{|\mathcal{D}|}\sum_{\mathbf{X}\in\mathcal{D}}(\frac{1}{T\cdot|\mathcal{Z}|}\sum_{t=1}^{T}\sum_{d\in\mathcal{Z}}\big{(}d(\mathbf{X},\tilde{\mathbf{X}})+1)^{-1},$
-where $\mathcal{Z}$ is the set of zero-valued dimensions in $\mathbf{X}$.
-By penalizing dimensions that deviate unrealistically in regions of zero value, this loss encourages the model to generate plausible anomalies while preserving the unique characteristics of zero-valued dimensions.
-3.2.5. Enhanced KL Divergence Loss
-The standard KL divergence loss is typically used to regularize the latent space by aligning the approximate posterior distribution $q(z|\mathbf{X})$, parameterized by $\mu$ and $\sigma^{2}$, with a standard Gaussian prior $p(z)$, defined as $\mathcal{N}(0,1)$.
-While effective for many tasks, this standard approach may fail to create the tight and compact latent representations necessary for generating anomalies that are distinguishable from normal samples.
-To address this, we propose an enhanced KL divergence loss by incorporating a tunable prior variance, $\sigma_{\text{prior}}^{2}$, where $\sigma_{\text{prior}}<1$.
-This modification enforces tighter representations of normal samples in the latent space.
-The modified KL divergence loss is defined as:
-(5) $\mathcal{L}_{\text{en-KL}}=-\frac{1}{2}\sum_{i=1}^{|\mathbf{\mathcal{D}}|}\sum_{j=1}^{L}\Big{[}1+\log(\sigma^{2}_{ij})-\mu_{ij}^{2}-\frac{\sigma^{2}_{ij}}{\sigma_{\text{prior}}^{2}}+2\log(\sigma_{\text{prior}})\Big{]},$
-Where $\mu_{ij}$ and $\sigma^{2}_{ij}$ are the mean and variance of the latent distribution for the $j$-th dimension of the $i$-th sample, and $\sigma_{\text{prior}}$ is the prior’s standard deviation.
-Theorem 1 (Enhanced Separation Between Normals and Anomalies in Latent Space).
+3 GenIAS Method
+Problem Definition.
+Let $\mathcal{D}$ be a time series dataset consisting of $N$ windows, where each window $\mathbf{X}\in\mathbb{R}^{T\times D}$ has length $T$ and $D$ variables, drawn from an unknown data distribution $p(\mathbf{X})$.
+In typical TSAD settings, $\mathcal{D}$ is dominated by normal samples, while anomalies are rare.
+The objective of anomaly generation is to learn a mapping $\Gamma:\mathbf{X}\rightarrow\widetilde{\mathbf{X}}$ that transforms a normal window $\mathbf{X}$ into an anomalous window $\widetilde{\mathbf{X}}\in\mathbb{R}^{T\times D}$.
+The generated anomalies should preserve the local temporal context of $\mathbf{X}$ while introducing controlled deviations that make $\widetilde{\mathbf{X}}$ distinguishable from normal patterns and informative for training TSAD models.
+For anomaly generation via latent space perturbation, we identify two key considerations for generating anomaly-discriminative samples.
+The first is to design a latent space perturbation mechanism for $\mathbf{z}\in\mathbb{R}^{L}$, parameterized by $p(\mathbf{z}\mid\mathbf{X})$, that injects controlled deviations into the latent representation, which serves as the input to the decoder for construct anomalous windows, to generate anomalies that satisfy the above properties.
+The second is to densify normal representations in latent space, which reduces overlap between normal samples and generated anomalies.
+This is applicable to both univariate and multivariate settings.
+3.1 Anomaly Generation using GenIAS
+GenIAS Overview.
+The core of GenIAS is a novel learnable perturbation mechanism operating on a variational latent space that is explicitly learned through a tailored training strategy, where distributions of normal time series data and the perturbation mechanism are jointly optimized.
+This joint optimization is motivated by the above design considerations, ensuring that the learned perturbations align well with the context of the normal time series distribution while being informative for anomaly generation.
+Generation Workflow.
+As shown in Figure 2, GenIAS adopts an encoder-decoder architecture, where a Temporal Convolutional Network (TCN) is used as the backbone to extract complex temporal patterns and dependencies through stacked convolutional layers interleaved with nonlinear activations and dropout.
+A variational autoencoder is then used to model a continuous latent Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$, where $\mu$ and $\sigma^{2}$ are $L$-dimensional vectors capturing the central tendency (mean) and spread (variance) of normal patterns, respectively, providing a basis for controlled anomaly generation.
+A normal window can be generated by first sampling a latent vector $\mathbf{z}=\mu+(\sigma\odot\epsilon)$, which is then passed to the decoder to reconstruct the corresponding time series window.
+Anomaly Injection Pipeline.
+GenIAS employs a two-stage pipeline for anomaly generation, consisting of anomalous window generation and deviation-based patching.
+Learnable perturbations are applied during latent sampling of a trained VAE to generate anomalous latent vectors (Eq.1), which are decoded into anomalous windows (detailed in Algo.
+1).
+While the reconstructed anomalous window captures coherent abnormal patterns, directly replacing the full window can be insufficient or overly aggressive for TSAD, as real anomalies are often localized.
+We therefore apply a deviation-based patching strategy that selectively replaces entire dimensions from the normal trajectory with their anomalous counterparts only when a non-trivial deviation is detected.
+The patching mechanism is detailed in Sec.3.3.
+Deviation via Variance Perturbation.
+Although VAEs are commonly used for modelling and generating normal data, we exploit their variational latent distributions to generate anomalies by adjusting the variance around the latent mean to produce desired deviations.
+We propose perturbing the variance of the learned VAE instead of the mean so that the generated anomalies follow the temporal property and central tendency of the times series.
+To generate anomalous representations $\tilde{\mathbf{z}}$, the standard deviation $\sigma$ is perturbed while keeping $\mu$ fixed:
+$\vskip-3.0pt\tilde{\mathbf{z}}=\mu+\psi\cdot(\sigma\odot\epsilon),\vskip-2.0pt$ (1)
+where $\psi$ is a learned perturbation scale, and $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is a noise vector.
+We find that perturbing the latent mean can introduce large deviations, but it violates the central tendency, often exhibiting out of context temporal properties that are less informative for representing realistic anomalies.
+3.2 GenIAS’s Joint Learning
+We jointly learn the perturbation scale and the time series latent distribution to ensure compatibility.
+In addition to the reconstruction loss for encoding the context of time series (Eq.
+5), we introduce two constraints to tighten the variance of the normal latent distribution and to regularize an effective perturbation scale.
+Specifically, we enforce a compact normal manifold through a compact loss (Eq.
+2) to reduce overlap between generated anomalies and normal data, and calibrate the perturbation scale to produce context-aligned deviations using a perturbation loss (Eq.
+3).
+Intuitively, effective synthetic anomalies should deviate from normal patterns without leaving the local context, since overly subtle deviations label near normal samples as anomalies and introduce label noise, while overly large deviations often become out of context and uninformative.
+Compact KL Divergence for Distinguishability.
+The standard KL divergence loss is mainly used to regularize the latent space by aligning the approximate posterior distribution $q(z|\mathbf{X})$, parameterized by $\mu$ and $\sigma^{2}$, with a standard Gaussian prior $p(z)$, defined as $\mathcal{N}(0,1)$.
+However, this does not explicitly control the spread of the normal latent distribution, which can become overly diffuse and increase overlap between generated anomalies and normal samples.
+We address this by introducing a compact KL divergence loss with a tunable prior variance $\sigma_{\text{prior}}^{2}$:
+$\mathcal{L}_{\text{comp-KL}}=-\frac{1}{2}\sum_{i=1}^{|\mathbf{\mathcal{D}}|}\sum_{j=1}^{L}\Big[1+\log(\sigma^{2}_{ij})-\mu_{ij}^{2}-\frac{\sigma^{2}_{ij}}{\sigma_{\text{prior}}^{2}}+2\log(\sigma_{\text{prior}})\Big],$ (2)
+Where $\mu_{ij}$ and $\sigma^{2}_{ij}$ are the mean and variance of the latent distribution for the $j$-th dimension of the $i$-th sample, and $\sigma_{\text{prior}}<1$ is the prior standard deviation.
+We show in Lemma 3.1 that introducing the tunable prior variance $\sigma_{\text{prior}}$ leads to tighter latent representations.
+Lemma 3.1 (Compact Latent Space in Compact KL Loss).
+Consider the KL divergence between the estimated posterior and the prior distributions for a latent dimension $j$.
+If the compact KL loss (Eq.
+2) is minimized (or its gradient dominates locally) with a prior variance $\sigma^{2}_{\text{prior}}<1$, then for each latent variable $z_{j}$ the optimal posterior variance satisfies $\sigma^{2}_{j}=\sigma^{2}_{\text{prior}}$.
+Consequently, the posterior variance of each latent variable is driven toward the prior variance $\sigma^{2}_{\text{prior}}$, thereby imposing a compact spread in the latent space.
+The full proof is provided in D.1.
+We now show in the following proposition that limiting the latent variance improves the distinguishability between generated anomalous and normal latent distributions under variance scaling perturbations.
+Proposition 3.2 (Compactness leads to greater separation in the latent space).
 Let a VAE be trained on normal data with a latent prior $\mathcal{N}(0,\sigma^{2}_{\text{prior}})$, where $\sigma_{\text{prior}}<1$.
@@
 If the encoder $\phi$ applies a perturbation scale $\psi$, inflating variance as $\sigma^{2}_{\text{anom}}=\psi\sigma^{2}_{\text{normal}},\quad\text{with }\psi>1,$ then the KL divergence between normal and anomalous latent distributions strictly increases compared to the case where $\sigma_{\text{prior}}\geq 1$.
-The proof is provided in Appendix D.3. We also illustrate in Figure 8 in the Appendix how our enhanced KL loss leads to the separation of reconstruction error in the raw space.
-The benefits of the enhanced KL divergence loss are as follows:
-  • Controlled spread: $\sigma_{\text{prior}}$ determines the spread of normal samples in the latent space.
-Smaller values enforce tighter clustering, leading to compact representations while allowing limited diversity for variability.
-(See Lemma 1)
-  • Regularization: The KL loss with $\sigma_{\text{prior}}$ ensures a well-regularized latent space, reducing overfitting and promoting smooth sampling.
-  • Improved normal sample representation: By constraining the latent space, normal samples occupy a distinct, dense region, enhancing their separability from anomalies. (See Lemma 2 and Theorem 1)
-  • Enhanced interpretability: The structured latent space shaped by $\sigma_{\text{prior}}$ provides consistent and interpretable embeddings for both normal and anomalous data.
-3.2.6. Total Loss
-The total loss integrates all the above components, each weighted by a corresponding hyperparameter:
-(6) $\mathcal{L}_{\text{total}}=\alpha\cdot\mathcal{L}_{\text{recon}}+\beta\cdot\mathcal{L}_{\text{perturb}}+\gamma\cdot\mathcal{L}_{\text{zero-perturb}}+\zeta\cdot\mathcal{L}_{\text{en-KL}},$
-Where $\alpha$, $\beta$, $\gamma$, and $\zeta$ are the weights for the reconstruction, perturbation, zero-perturbation, and KL divergence losses, respectively.
-3.3. Deviation Patching Approach
-After training the GenIAS model in the first step, the anomaly injection process uses the pretrained model to perturb the latent space representation of time series windows ($\mathbf{X}$) and generate realistic anomalous samples ($\tilde{\mathbf{X}}$).
-This step consists of two phases: (1) anomaly injection via latent space perturbation and (2) patching the injected anomalies using a deviation-based approach, defined by a scaling factor $\tau$, to ensure the generated samples are realistic and aligned with the domain requirements.
-The injection and patching process is applied independently for each dimension ($d$) of the MTS, and more details are provided in Algorithm 2.
-Figure 3. Effect of deviation-based patching with different thresholds (0.4, 0.2, 0.05) on generated anomalous windows across various entities from the MSL, SMAP, SMD, Yahoo, and KPI datasets.
-4. Experiments
-We conduct extensive evaluations of GenIAS’s performance using a comprehensive selection of datasets and baselines.
-Our evaluation consists of two parts: (1) Generation Evaluation, where we compare GenIAS against other anomaly injection/generation models to assess its capability in generating realistic and diverse anomalies, and (2) Case Study – TSAD Evaluation, where we benchmark GenIAS against seventeen competing methods using nine real-world datasets covering both MTS and UTS.
-Specifically, to study the effectiveness of GenIAS’s anomaly generation in TSAD, we integrate GenIAS with the anomaly detector architecture from CARLA to implement a GenIAS-enhanced TSAD model (i.e., substituting CARLA’s anomaly generator with GenIAS while keeping the rest of the framework identical).
-We choose CARLA as our base detector due to its SOTA performance, which enables direct comparison with the strongest generation-based TSAD method.
-Through this comprehensive evaluation, we provide a thorough analysis of GenIAS’s capabilities in both anomaly generation and TSAD tasks.
-Table 1. An overview of dataset statistics.
-Benchmark # datasets # dims Train size Test size ARMSL (Hundman et al., 2018) 27 55 58,317 73,729 10.72% SMAP (Hundman et al., 2018) 55 25 140,825 444,035 13.13% SMD (Su et al., 2019) 28 38 708,405 708,420 4.16% SWaT (Mathur and Tippenhauer, 2016) 1 51 496,800 449,919 12.33% GECCO (Moritz et al., 2018) 1 9 69,260 69,261 1.1% SWAN (Angryk et al., 2020) 1 38 60,000 60,000 32.6% UCR (Wu and Keogh, 2021) 205 1 2,238,349 6,143,541 0.6% Yahoo-A1 (Laptev et al., 2015) 67 1 46,667 46,717 1.76% KPI (aio, 2018) 29 1 1,048,576 2,918,847 1.87%
-4.1. Benchmark Datasets
-GenIAS is thoroughly evaluated on nine widely adopted real-world benchmark datasets for TSAD.
-Their key statistics are reported in Table 1.
-In particular, our selection includes six MTS datasets (MSL, SMAP, SMD, SWaT, GECCO, and SWAN) and three UTS datasets (UCR, Yahool-A1 and KPI), containing diverse anomaly types and a variety of anomaly ratios, thereby providing a comprehensive evaluation of the performance of our proposed methods and the baselines.
-Detailed dataset descriptions can be found in Appendix A.
-4.2. Baselines
-GenIAS is compared with seventeen relevant and competitive baseline methods that leverage a wide array of anomaly detection approaches, which can be categorised into four broad categories: classic anomaly detection methods, positive-unlabeled (P-U) methods, unsupervised deep TSAD methods that do not leverage anomaly generation (US-WI), and unsupervised TSAD methods that employ anomaly generation (US-I).
-Specifically, three classic general AD methods - OCSVM (Schölkopf et al., 1999), LOF (Breunig et al., 2000), and Isolation Forest (Fei Tony Liu and Zhou, 2008) - are included to establish a base performance for demonstrating the advantages of deep learning TSAD approaches.
-LSTM-VAE (Park et al., 2018), a deep learning TSAD method that utilises labeled anomalies, is included to examine the effects of leveraging labelled anomalies for TSAD.
-US-WI is another relevant category, which includes a variety of unsupervised TSAD baselines that do not leverage anomaly generation.
-Donut (Xu et al., 2018), OmniAnomaly (Su et al., 2019), AnomalyTransformer (Xu et al., 2021), and TranAD (Tuli et al., 2022b) use reconstruction error to detect anomalies.
-THOC (Shen et al., 2020b) and TimesNet (Wu et al., 2023) treat TSAD as a forecasting problem.
-MTAD-GAT (Zhao et al., 2020) integrates graph attention mechanisms to capture both temporal and multivariate dependencies.
-Representation learning methods like TS2Vec (Yue et al., 2022) and DCdetector (Yang et al., 2023) focus on establishing robust and anomaly-discriminative representations.
-US-I is the most relevant category, which allows us to compare GenIAS with existing generation methods for unsupervised TSAD.
-We include four recent unsupervised TSAD methods that leverage various anomaly generation approaches: NCAD (Carmona et al., 2022), CutAddPaste (Wang et al., 2024), COUTA (Xu et al., 2024), and CARLA (Darban et al., 2025).
-4.3. Evaluation Metrics
-We adopt seven evaluation metrics to assess GenIAS’s performance in terms of generation quality and TSAD performance.
+The proof is provided in D.3.
+The increased KL divergence under lower variance implies statistically improved distinguishability.
+We also illustrate in Fig.
+8 in the Appendix how our compact KL loss leads to the separation of reconstruction error in the raw space.
+Perturbation Loss for Informative Deviation. We define a perturbation loss to control the magnitude of deviation introduced by latent perturbations composed of two complementary terms.
+The first term is a triplet-style margin loss that enforces a minimum separation between the original sample $\mathbf{X}$ and the generated anomalous sample $\widetilde{\mathbf{X}}$, relative to the reconstructed normal sample $\hat{\mathbf{X}}$.
+This term ensures that generated anomalies deviate sufficiently from the original time series by pushing anomalous windows away from normal windows in the raw time series space with a minimum margin $\delta_{\text{min}}$.
+[…]
+$\mathcal{L}_{\text{perturb}}=\;$ $\frac{1}{|\mathcal{D}|}\sum_{\mathbf{X}\in\mathcal{D}}\Big[\max\big(d(\mathbf{X},\hat{\mathbf{X}})-d(\mathbf{X},\tilde{\mathbf{X}})+\delta_{\text{min}},0\big)\Big]$
+$+\frac{1}{|\mathcal{D}|}\sum_{\mathbf{X}\in\mathcal{D}}\max\big(d(\mathbf{X},\tilde{\mathbf{X}})-\delta_{\text{max}},0\big)$ (3)
+These terms constrain the perturbation magnitude to avoid overly obvious deviations while ensuring anomalies remain distinguishable.
+Zero-Perturbation Loss for Zero-valued Dimensions.
+Time series windows may contain dimensions that are entirely zero, corresponding to inactive or sparse signals.
+Directly applying perturbations to such dimensions can result in unrealistic anomalies.
+We introduce a zero-perturbation loss that regularizes deviations in zero-valued dimensions.
+Specifically, the loss penalizes large deviations in dimensions where the original signal is zero, encouraging perturbations to remain small and plausible while preserving the structural characteristics of these dimensions.
+The loss is defined as:
+$\mathcal{L}_{\text{zero-perturb}}=\frac{1}{|\mathcal{D}|}\sum_{\mathbf{X}\in\mathcal{D}}\big(\frac{1}{T\cdot|\mathcal{Z}|}\sum_{t=1}^{T}\sum_{d\in\mathcal{Z}}\big(d(\mathbf{X},\tilde{\mathbf{X}})+1)^{-1}\big),$ (4)
+where $\mathcal{Z}$ denotes the set of zero-valued dimensions in $\mathbf{X}$.
+This term discourages unrealistic perturbations in zero-valued channels while allowing controlled deviations when appropriate.
+Reconstruction Loss for Learning Normal.
+The reconstruction loss is widely used for learning latent representation and is defined as Mean Squared Error (MSE) between an input sample $\mathbf{X}$ and its reconstruction $\hat{\mathbf{X}}$:
+$\mathcal{L}_{\text{recon}}=\frac{1}{|\mathcal{D}|}\sum_{\mathbf{X}\in\mathcal{D}}\frac{1}{T\cdot D}\sum_{t=1}^{T}\sum_{d=1}^{D}\big(\mathbf{X}[t,d]-\hat{\mathbf{X}}[t,d]\big)^{2},$ (5)
+which is included to capture the temporal properties of normal samples as the basis for perturbation.
+Overall Objective. The overall objective for GenIAS joint learning can be defined as:
+$\mathcal{L}_{\text{total}}=\alpha\cdot\mathcal{L}_{\text{recon}}+\beta\cdot\mathcal{L}_{\text{perturb}}+\gamma\cdot\mathcal{L}_{\text{zero-perturb}}+\zeta\cdot\mathcal{L}_{\text{comp-KL}},$ (6)
+where $\alpha$, $\beta$, $\gamma$, and $\zeta$ are the coefficients for the reconstruction, perturbation, zero-perturbation, and KL divergence losses, respectively.
+Figure 3: Effect of deviation-based patching thresholds on generated anomalies across SMAP and Yahoo datasets.
+3.3 Deviation Patching Approach
+As discussed in Sec.3.1, directly using reconstructed anomalous windows has practical limitations.
+We therefore introduce a deviation-based patching strategy to selectively inject anomalous components while preserving normal context.
+Specifically, patching is performed independently for each dimension $d$ according to $\mathbf{X}_{\text{patched},d}=\widetilde{\mathbf{X}}_{d}$ if $\lVert\mathbf{X}_{d}-\widetilde{\mathbf{X}}_{d}\rVert^{2}>\tau\cdot(\max(\mathbf{X}_{d})-\min(\mathbf{X}_{d}))$, otherwise $\mathbf{X}_{d}$, where $\tau$ is a scaling factor controlling the sensitivity of anomaly injection.
+This strategy replaces entire variable trajectories only when a non-trivial deviation is detected, preserving normal patterns in unaffected dimensions.
+We show its effect in Fig.
+3.
+The complete procedure is summarized in Algo 2.
+4 Experiments
+We extensively evaluate GenIAS using a diverse set of datasets and baselines in two parts: (1) Generation Evaluation, comparing GenIAS with other anomaly injection models to assess its capability to produce realistic and diverse anomalies; and (2) Case Study – TSAD Evaluation, benchmarking GenIAS against seventeen competing methods using nine real-world MTS and UTS datasets.
+To assess its impact on TSAD, we replace CARLA’s anomaly generator with GenIAS while keeping the rest of the architecture unchanged.
+CARLA is selected as the base detector for its SOTA performance, enabling fair comparison with the strongest injection-based method.
+This comprehensive evaluation allows for a detailed analysis of GenIAS’s effectiveness in both anomaly generation and TSAD.
+Table 1: An overview of dataset statistics.
+Benchmark # datasets # dims Train size Test size AR% MSL (Hundman et al., 2018) 27 55 58,317 73,729 10.72% SMAP (Hundman et al., 2018) 55 25 140,825 444,035 13.13% SMD (Su et al., 2019) 28 38 708,405 708,420 4.16% SWaT (Mathur and Tippenhauer, 2016) 1 51 496,800 449,919 12.33% GECCO (Moritz et al., 2018) 1 9 69,260 69,261 1.1% SWAN (Angryk et al., 2020) 1 38 60,000 60,000 32.6% UCR (Wu and Keogh, 2021) 250 1 2,238,349 6,143,541 0.6% Yahoo-A1 (Laptev et al., 2015) 67 1 46,667 46,717 1.76% KPI (1) 29 1 1,048,576 2,918,847 1.87%
+4.1 Benchmark Datasets
+GenIAS is thoroughly evaluated on nine widely used real-world TSAD benchmark datasets, with key statistics reported in Table 1.
+The selection includes six MTS and three UTS datasets, covering diverse anomaly types and ratios to ensure a comprehensive comparison against baselines.
+Detailed dataset descriptions are provided in Appendix A.
+4.2 Baselines
+GenIAS is compared with seventeen competitive baseline methods that leverage a wide range of anomaly detection models, which can be categorized into four broad categories: classic anomaly detection methods, positive-unlabeled (P-U) methods, unsupervised deep TSAD methods that do not leverage anomaly generation (US-WI), and unsupervised TSAD methods that employ anomaly generation (US-I).
+Specifically, three classic general AD methods— OCSVM (Schölkopf et al., 1999), LOF (Breunig et al., 2000), and Isolation Forest (Fei Tony Liu and Zhou, 2008)—to demonstrate the advantages of deep learning TSAD approaches.
+LSTM-VAE (Park et al., 2018) is included to assess the impact of incorporating labeled data in TSAD.
+US-WI includes a range of unsupervised TSAD baselines that do not use anomaly generation.
+Methods such as Donut (Xu et al., 2018), OmniAnomaly (Su et al., 2019), AnomalyTransformer (Xu et al., 2021), and TranAD (Tuli et al., 2022a) rely on reconstruction error for anomaly detection.
+THOC (Shen et al., 2020b) and TimesNet (Wu et al., 2023) approach TSAD as a forecasting task, while MTAD-GAT (Zhao et al., 2020) incorporates graph attention to model temporal and multivariate dependencies.
+Representation learning methods like TS2Vec (Yue et al., 2022) and DCdetector (Yang et al., 2023) aim to learn robust, anomaly-discriminative embeddings.
+US-I, the most relevant category for GenIAS, includes unsupervised TSAD methods that use anomaly generation.
+We compare GenIAS with four recent models in this category: NCAD (Carmona et al., 2022), CutAddPaste (Wang et al., 2024), COUTA (Xu et al., 2024), and CARLA (Darban et al., 2025).
+4.3 Evaluation Metrics
+We adopt seven metrics to evaluate GenIAS’s performance in terms of generation quality and TSAD performance.
 Generation quality.
-To quantitatively examine the quality of anomaly generation, we introduce two key metrics: Anomalous Representation Proximity (ARP), which assesses realism, and Entropy-Based Diversity Index (EDI), which characterizes the diversity of the generated anomalies, motivated by the idea in AnomalyDiffusion (Hu et al., 2024).
-Due to page limits, the definitions and discussions of these metrics are provided in the Appendix E.
-In general, ARP is the primary metric as it measures the closeness of the generated anomalies to the actual anomalies, see Appendix E.1.
-EDI is complementary to ARP, as diversity should only be examined under the condition that the generated anomalies can illustrate the actual anomalies, see Appendix E.2.
+To quantitatively examine anomaly generation quality, we introduce two key metrics: Anomalous Representation Proximity (ARP), which assesses realism, and Entropy-Based Diversity Index (EDI), which characterizes the diversity, motivated by the idea in AnomalyDiffusion (Hu et al., 2024).
+Due to space constraints, definition and further discussion are in Appendix E.
+In general, ARP is the primary metric, reflecting how closely generated anomalies resemble actual ones, while EDI is complementary to ARP by assessing diversity only when realism is satisfied.
 TSAD evaluation.
-We consider five widely-used, complementary metrics: Best F1 score, AUPR (Area Under the Precision-Recall Curve), Affiliation F1  (Huet et al., 2022), PATE  (Ghorbani et al., 2024), and AUROC (Area Under the Receiver Operating Characteristic Curve) for TSAD performance evaluation.
-This combination of metrics provides a comprehensive, fair and robust TSAD evaluation.
-Detailed descriptions of evaluation metrics are provided in Appendix E.3.
-4.4. Implementation Details
-Our experiments are conducted on a server with an NVIDIA A40 GPU, 13 CPU cores, and 250GB RAM.
-We employ the unsupervised TSAD network architecture and training scheme from CARLA as our default TSAD model to perform GenIAS-enhanced detection.
+We consider five widely-used, complementary metrics to evaluate TSAD performance: F1 score, AUPR (Area Under the Precision-Recall Curve), Affiliation F1  (Huet et al., 2022), PATE  (Ghorbani et al., 2024), and AUROC (Area Under the Receiver Operating Characteristic Curve).
+A combination of metrics offers a comprehensive, fair, and robust assessment.
+See Appendix E.3 for details.
+4.4 Implementation Details
+Experiments are run on a server with an NVIDIA A40 GPU, 13 CPU cores, and 250GB RAM.
+We adopt the unsupervised TSAD architecture and training scheme from CARLA as the default model for GenIAS-enhanced detection.
 GenIAS is implemented in Python using PyTorch and trained using the Adam optimizer with a learning rate of $10^{-4}$ for 1000 epochs.
-The default window size $T$ is set to 200, and latent dimension size $L$ is set to 100 and 50, for MTS and UTS respectively.
+The default window size $T$ is set to 200, and the latent dimension size $L$ is set to 100 and 50, for MTS and UTS, respectively.
 The hyperparameters for the total loss are $\alpha=1.0$, $\beta=0.1$, $\gamma=0.0/0.01$ (for UTS/MTS), and $\zeta=0.1$, which are fixed for all datasets.
-We adopt the baseline results that are produced using the original implementation and maintain their default hyperparameter settings.
-Due to the page limit, the detailed implementation details and hyperparameter settings are provided in Appendix F.
-Table 2. Anomaly generation quality of GenIAS and US-I baselines in ARP, with the best model in bold and the second-best model underlined, where CAP abbreviates CutAddPaste.
+We use baseline results from their original implementations with default hyperparameters.
+Detailed implementation and hyperparameter settings are provided in Appendix F.
+4.5 Generation Performance
+Table 2: Evaluation results of anomaly generation quality for GenIAS and US-I baselines on the ARP metric, with the best in bold and second-best underlined (CAP: CutAddPaste).
 Metric Dataset COUTA NCAD CAP CARLA GenIAS ARP MSL 0.3613 0.3568 0.3406 0.3586 0.3605 SMAP 0.7573 0.8052 0.8110 0.7636 0.8113 SMD 0.0951 0.0810 0.0896 0.0958 0.0922 SWaT 0.9970 0.9964 0.9963 0.9980 0.9974 GECCO 0.1296 0.1145 0.1335 0.1163 0.1265 SWAN 0.2400 0.2362 0.2313 0.2410 0.2415 Yahoo 0.6541 0.6139 0.6476 0.6632 0.6756 KPI 0.6856 0.8427 0.7207 0.7582 0.7907 UCR 0.1998 0.2059 0.1936 0.2086 0.2161 Rank 3.11 3.78 3.89 2.44 1.78 EDI MSL 0.8251 0.7311 0.7169 0.5822 0.8833 SMAP 0.8557 0.6699 0.6364 0.6335 0.8932 SMD 0.8782 0.7648 0.6404 0.8041 0.9262 SWaT 0.6425 0.6164 0.7967 0.6005 0.7259 GECCO 0.7379 0.5571 0.7072 0.9327 0.8303 SWAN 0.9589 0.9189 0.7428 0.9286 0.9840 Yahoo 0.8627 0.7202 0.7577 0.5948 0.9182 KPI 0.7607 0.6166 0.7505 0.6962 0.8442 UCR 0.8166 0.7321 0.7536 0.7483 0.8961 Rank 2.22 4.11 3.56 3.89 1.22
-4.5. Generation Performance
-We first examine the anomaly generation quality of GenIAS and the competing methods that employ anomaly injection (i.e., the US-I category, see section 4.2) in terms of ARP and EDI, as discussed in Section 4.3.
+We first examine the anomaly generation quality of GenIAS and the competing methods that employ anomaly injection (i.e., the US-I category, see section 4.2) in terms of ARP and EDI (Section 4.3).
 We report the results of ARP and EDI in Table 2.
-GenIAS achieves the highest average ARP ranking, suggesting it generates overall the most realistic anomalies across all datasets.
-The realism of the anomaly distribution is essential for providing anomaly-informed supervision in the absence of labelled anomalies.
-Similarly for EDI, GenIAS also achieves the highest average ranking, indicating good diversity in its generated anomalies while sufficiently illustrating the actual anomalies.
-This diversity implies that the generated anomalies can also lie beyond the original anomaly distribution, which can be beneficial for learning a more stringent normal class boundary, further improving TSAD performance, provided that they do not interfere with the identification of the normal region.
+GenIAS achieves the highest average ARP ranking, suggesting it generates, overall, the most realistic anomalies across all datasets.
+The realism of the anomaly distribution is essential for providing anomaly-informed supervision in the absence of labeled anomalies.
+Similarly, for EDI, GenIAS also achieves the highest average ranking, indicating good diversity in its generated anomalies while sufficiently illustrating the actual anomalies.
+This diversity implies that the generated anomalies can also lie beyond the original anomaly distribution, which can be beneficial for learning a stricter normal class boundary, further improving TSAD performance, provided that they do not interfere with the identification of the normal region.
 To achieve this, GenIAS treats creating informative but not excessive diversity as a key consideration through its novel perturbation loss.
-Table 3.
-TSAD performance of GenIAS and baseline methods using F1 score and AUPR, with the best model in bold and the second-best model underlined.
+4.6 Visualization
+Figure 4: Time series injection methods on the Yahoo dataset. Real anomalies are shown in blue, and synthetic anomalies in orange.
+Table 3: TSAD performance of GenIAS and baseline methods using F1 score, with the best model in bold and the second-best model underlined.
 U and M denote results unavailable for models limited to UTS or MTS, respectively.
-MTS UTS Metric Cat.
-Methods MSL SMAP SMD SWAT GECCO SWAN Avg-M Yahoo KPI UCR Avg-U Avg F1 Classic OCSVM (1999) 0.308 0.411 0.319 0.743 0.297 0.557 0.439 0.608 0.148 0.011 0.256 0.378 LOF (2000) 0.368 0.400 0.204 0.331 0.075 0.655 0.339 0.586 0.101 0.010 0.232 0.303 IForest (2008) 0.278 0.353 0.306 0.745 0.255 0.717 0.442 0.553 0.236 0.010 0.266 0.384 PU LSTM-VAE (2018) 0.407 0.437 0.298 0.724 0.208 0.625 0.450 M M M M M US-WI Donut (2018) U U U U U U U 0.489 0.126 0.009 0.208 U Omni (2019) 0.243 0.325 0.459 0.763 0.166 0.541 0.416 M M M M M THOC (2020) 0.309 0.327 0.168 0.638 0.146 0.506 0.349 0.253 0.233 0.010 0.165 0.288 MTAD-GAT(2020) 0.473 0.519 0.347 0.242 0.313 0.599 0.416 M M M M M AnomalyTran (2021) 0.345 0.407 0.304 0.738 0.155 0.524 0.412 M M M M M TranAD (2022) 0.428 0.472 0.361 0.310 0.280 0.527 0.396 0.566 0.287 0.010 0.288 0.360 TS2Vec (2022) 0.299 0.371 0.173 0.261 0.063 0.502 0.278 0.484 0.204 0.007 0.232 0.263 TimesNet (2023) 0.358 0.402 0.338 0.217 0.297 0.492 0.351 0.513 0.241 0.011 0.255 0.319 Dcdetector (2023) 0.227 0.275 0.083 0.217 0.021 0.492 0.219 0.112 0.042 0.008 0.054 0.164 US-I NCAD (2022) 0.266 0.358 0.183 0.217 0.296 0.492 0.302 0.134 0.168 0.020 0.107 0.237 CutAddPaste (2024) 0.363 0.458 0.192 0.700 0.280 0.896 0.482 0.321 0.620 0.042 0.328 0.430 COUTA (2024) 0.414 0.421 0.362 0.312 0.389 0.626 0.421 0.658 0.221 0.010 0.296 0.379 CARLA (2025) 0.523 0.529 0.511 0.721 0.293 0.908 0.581 0.749 0.308 0.039 0.365 0.509 Ours GenIAS 0.579 0.620 0.528 0.755 0.329 0.937 0.625 0.811 0.393 0.064 0.423 0.557 AUPR Classic OCSVM (1999) 0.202 0.256 0.327 0.696 0.163 0.608 0.375 0.703 0.300 0.028 0.344 0.365 LOF (2000) 0.217 0.248 0.277 0.185 0.031 0.624 0.264 0.615 0.109 0.043 0.256 0.261 IForest (2008) 0.182 0.190 0.278 0.739 0.162 0.739 0.382 0.614 0.288 0.021 0.308 0.357 PU LSTM-VAE (2018) 0.285 0.258 0.395 0.685 0.045 0.674 0.390 M M M M M US-WI Donut (2018) U U U U U U U 0.264 0.078 0.015 0.119 U Omni (2019) 0.149 0.115 0.365 0.713 0.034 0.587 0.327 M M M M M THOC (2020) 0.240 0.195 0.107 0.537 0.076 0.432 0.265 0.349 0.229 0.014 0.197 0.242 MTAD-GAT(2020) 0.335 0.339 0.401 0.095 0.200 0.613 0.331 M M M M M AnomalyTran (2021) 0.236 0.264 0.273 0.680 0.059 0.568 0.347 M M M M M TranAD (2022) 0.273 0.287 0.412 0.192 0.119 0.586 0.312 0.691 0.285 0.022 0.333 0.319 TS2Vec (2022) 0.132 0.148 0.113 0.136 0.040 0.408 0.163 0.491 0.221 0.009 0.240 0.189 TimesNet (2023) 0.283 0.208 0.385 0.082 0.189 0.508 0.276 0.671 0.237 0.032 0.313 0.288 Dcdetector (2023) 0.129 0.124 0.043 0.126 0.011 0.326 0.127 0.041 0.018 0.009 0.023 0.092 US-I NCAD (2022) 0.146 0.156 0.096 0.106 0.135 0.407 0.174 0.067 0.089 0.103 0.086 0.145 CutAddPaste (2024) 0.242 0.322 0.153 0.582 0.159 0.808 0.378 0.116 0.602 0.188 0.302 0.352 COUTA (2024) 0.247 0.248 0.400 0.159 0.264 0.685 0.334 0.555 0.282 0.049 0.295 0.321 CARLA (2025) 0.501 0.448 0.507 0.681 0.201 0.814 0.525 0.627 0.299 0.247 0.391 0.481 Ours GenIAS 0.529 0.467 0.512 0.704 0.236 0.882 0.555 0.794 0.357 0.280 0.477 0.529
-4.6. Visualization
-(a) Yahoo - Entity real_65
-(b) SMAP - Entity E5
-Figure 4.
-Each plot visualizes a different time series (TS) injection or generation method using t-SNE.
-Blue markers represent the original anomalous TS, while orange markers represent the injected or generated anomalous TS.
-To illustrate GenIAS’s enhanced anomaly generation compared to baseline methods that also leverage anomaly generation (US-I), we present t-SNE visualizations comparing their generated and real anomalies in Figure 4 on two selected datasets, Yahoo (UTS) and SMAP (MTS).
-Some additional visualizations are included in Figure 10 in Appendix G.2.3.
-Impressively, the distributions of the generated anomalies by GenIAS align with the actual anomalies distribution substantially closer compared to the baseline methods (i.e., actual anomaly regions are densely covered by the generated anomalies).
+MTS UTS Cat.
+Methods MSL SMAP SMD SWAT GECCO SWAN Avg-M Yahoo KPI UCR Avg-U Avg Classic OCSVM (1999) 0.308 0.411 0.319 0.743 0.297 0.557 0.439 0.608 0.148 0.011 0.256 0.378 LOF (2000) 0.368 0.400 0.204 0.331 0.075 0.655 0.339 0.586 0.101 0.010 0.232 0.303 IForest (2008) 0.278 0.353 0.306 0.745 0.255 0.717 0.442 0.553 0.236 0.010 0.266 0.384 PU LSTM-VAE (2018) 0.407 0.437 0.298 0.724 0.208 0.625 0.450 M M M M M US-WI Donut (2018) U U U U U U U 0.489 0.126 0.009 0.208 U Omni (2019) 0.243 0.325 0.459 0.763 0.166 0.541 0.416 M M M M M THOC (2020) 0.309 0.327 0.168 0.638 0.146 0.506 0.349 0.253 0.233 0.010 0.165 0.288 MTAD-GAT (2020) 0.473 0.519 0.347 0.242 0.313 0.599 0.416 M M M M M AnomalyTran (2021) 0.345 0.407 0.304 0.738 0.155 0.524 0.412 M M M M M TranAD (2022) 0.428 0.472 0.361 0.310 0.280 0.527 0.396 0.566 0.287 0.010 0.288 0.360 TS2Vec (2022) 0.299 0.371 0.173 0.261 0.063 0.502 0.278 0.484 0.204 0.007 0.232 0.263 TimesNet (2023) 0.358 0.402 0.338 0.217 0.297 0.492 0.351 0.513 0.241 0.011 0.255 0.319 Dcdetector (2023) 0.227 0.275 0.083 0.217 0.021 0.492 0.219 0.112 0.042 0.008 0.054 0.164 US-I NCAD (2022) 0.266 0.358 0.183 0.217 0.296 0.492 0.302 0.134 0.168 0.020 0.107 0.237 CutAddPaste (2024) 0.363 0.458 0.192 0.700 0.280 0.896 0.482 0.321 0.620 0.042 0.328 0.430 COUTA (2024) 0.414 0.421 0.362 0.312 0.389 0.626 0.421 0.658 0.221 0.010 0.296 0.379 CARLA (2025) 0.523 0.529 0.511 0.721 0.293 0.908 0.581 0.749 0.308 0.039 0.365 0.509 Ours GenIAS+CARLA 0.579 0.620 0.528 0.755 0.329 0.937 0.625 0.811 0.393 0.064 0.423 0.557
+To illustrate GenIAS’s enhanced anomaly generation compared to other anomaly injection-based baselines (US-I), we present t-SNE visualizations in Figure 4, comparing generated and real anomalies on Yahoo.
+Additional visualizations are provided in Figure 10 in Appendix G.2.3.
+Impressively, GenIAS-generated anomalies align substantially more closely with the actual anomaly distribution compared to baseline methods, effectively covering true anomaly regions.
 This is essential to provide anomaly-discriminative information for establishing more distinctive normal and anomaly representations in unsupervised TSAD training.
-On the other hand, the baseline methods, to varying degrees, are unable to generate anomalies that sufficiently represent the actual anomaly distribution, which is evidenced by major blue areas (representing actual anomaly regions) that are not covered by orange areas (generated anomaly regions).
-For example, CARLA, the second-best method for overall ARP and TSAD performance, also suffers from these issues: on Yahoo (a), generated anomalies largely miss the anomaly regions, while on SMAP (b), generated anomalies form two major clusters leaving a gap in between, which is an actual anomaly region failing to be represented.
-These observations further substantiate that relying on handcrafted, predefined anomaly patterns for anomaly generation is prone to distribution misalignment when the assumed patterns deviate from actual ones, and disconnected/scattered anomaly regions when the assumed patterns differ significantly.
-Whereas GenIAS, through controlled and regularized perturbation on the learned distribution, generates anomalies without pattern speculation, achieving both better alignment and smoothness.
-4.7. Main TSAD Performance
-We report the TSAD performance of GenIAS and the competing methods in terms of the best F1 score and AUPR in Table 3.
-Due to space limitations, please refer to the Appendix G.2.2 for results on the other three previously mentioned metrics discussed in Section 4.3.
-Impressively, GenIAS achieves significantly better overall TSAD performance compared to the baseline methods across all metrics, demonstrating its superior effectiveness in handling both MTS and UTS.
-Specifically, among methods that can operate on both types of time series, GenIAS achieves 9.4% and 29.5% performance improvements compared to the second and third best performing methods in F1, respectively.
-Similarly, it outperforms the second best model by 10.0% and third best model by 44.9% in terms of AUPR.
-Upon examining the more detailed TSAD performance with respect to factors such as method category and dataset type, we noted the following observations.
+In contrast, baseline methods struggle to generate anomalies that adequately capture the actual anomaly distribution, as shown by large blue regions (actual anomalies) left uncovered by orange regions (generated anomalies).
+For example, CARLA, the second-best method in overall ARP and TSAD performance, struggles on Yahoo, where its generated anomalies largely miss true anomaly regions.
+This highlights a key limitation of relying on handcrafted, predefined anomaly patterns: they risk distribution misalignment when assumptions deviate from reality, often resulting in disconnected or scattered anomaly regions.
+In contrast, GenIAS leverages controlled, regularized perturbation on the learned distribution to generate anomalies without pattern speculation, achieving better alignment and smoothness.
+4.7 Main TSAD Performance
+We report the TSAD performance of GenIAS and the competing methods in terms of the best F1 score in Table 3.
+Due to space constraints, results for the other four metrics discussed in Section 4.3 are provided in Appendix G.2.2.
+GenIAS achieves significantly better overall TSAD performance compared to the baselines across metrics, demonstrating its superior effectiveness on both MTS and UTS.
+Specifically, among methods applicable to both types of time series, GenIAS achieves F1 improvements of 9.4% and 29.5% over the second- and third-best, respectively.
+A more detailed analysis by method category and dataset type reveals the following key observations.
 Consistent effectiveness for both MTS and UTS.
 On each type of dataset, GenIAS yields similar levels of performance improvement as observed in overall performance.
-For example, compared to the second-best model for both types, GenIAS improves F1 by 7.6% and AUPR by 8.4% on MTS, and improves F1 by 15.9% and AUPR by 22.0% for UTS.
+For example, compared to the second-best model for both types, GenIAS improves F1 by 7.6% on MTS and improves F1 by 15.9%.
 This highlights its ability to generate informative and diverse anomalies that can mimic anomalous patterns within each dimension and the dependencies between dimensions.
 We also provide the average rank of all models in the form of critical difference diagrams (Demšar, 2006) for F1 and AUPR in both MTS and UTS (see Appendix G.2.1).
-GenIAS achieved the highest rank in all four comparisons and demonstrated statistically significant improvements over all baselines in MTS for both AUPR and F1.
+GenIAS achieved the highest rank in all four comparisons and demonstrated statistically significant improvements over all baselines in MTS for the F1 score.
 Enhanced Anomaly Generation.
-It is worth noting that even when compared to US-I baselines, GenIAS consistently outperforms them due to its ability to generate more realistic and diverse anomalous patterns.
-In particular, since GenIAS-enhanced TSAD model and CARLA only differ in their anomaly generation while sharing the same architecture and training methods, their performance gap underscores the effectiveness of GenIAS’s approach.
-This is attributed to its TSAD tailored variational representation learning and novel perturbation strategies, which enforce more distinctive normal representations and are not confined to predefined, handcrafted anomaly prototypes.
+GenIAS consistently outperforms US-I baselines by generating more realistic and diverse anomalous patterns.
+In particular, its performance gain over CARLA while only differing in their anomaly generation highlights the effectiveness of GenIAS’s tailored variational learning and novel perturbation strategies, which enforce more distinctive normal representations without relying on handcrafted anomaly types.
 Importance of anomaly generation in TSAD.
-It is notable that US-I injection methods are generally more effective than methods in other categories that do not leverage labeled anomalies.
-Remarkably, some of these methods, such as CutAddPaste and CARLA, can even outperform the supervised baseline LASM-VAE.
-This suggests that, with well-designed anomaly injection techniques, the generated pseudo anomalies can be sufficiently realistic to represent actual anomalies, including those not well captured by the labeled training data in supervised settings.
-Moreover, leveraging anomaly generation can generate both larger numbers and more diverse types of anomalies, which can improve TSAD representation learning.
-Therefore, anomaly injection should be incorporated as an essential module in TSAD models.
-4.8. Ablation Study
+US-I injection methods generally outperform other unsupervised approaches and, in some cases, like CutAddPaste and CARLA, can even outperform the supervised baseline LSTM-VAE.
+This suggests that well-designed anomaly generation can produce realistic pseudo-anomalies that capture true anomalies, including those missed by labeled data.
+Moreover, it enables greater anomaly diversity, improving representation learning.
+Therefore, anomaly injection should be considered as an essential module in TSAD models.
+4.8 Ablation Study
 To better understand the contributions of different components and configurations in our proposed model, we conducted an ablation study.
-Our analysis focuses on the following aspects: (i) The impact of the patch selection, (ii) The effectiveness of the prior variance, and (iii) The effectiveness of the min-max margin.
-4.8.1. Impact of the Patch Selection
-First, in Figure 5, we present our study of F1 and AUPR for two patching methods—Length-driven and Deviation patching—evaluated across varying portions of patch lengths and thresholds.
-The x-axis represents the portion of patch length (ranging from 0.2 to 1, in intervals of 0.2), while thresholds $\tau$ are explored at 0.05, 0.1, 0.2, 0.4, and 0.6.
-The best results for each method, identified through this parameter exploration, are then compared in Table 4.
-The results in Table 4 demonstrate that Deviation patching consistently outperforms the other approaches, achieving the highest F1 and AUPR scores across all datasets.
-For example, it achieves a 0.044 increase in F1 and a 0.063 improvement in AUPR for SMAP compared to the baseline without patching (wo Patch).
-Length-driven patching, which introduces anomalies based on predefined temporal lengths, shows moderate improvements over the baseline in most scenarios but slightly underperforms on more complex datasets, such as Yahoo, where it reduces F1 and AUPR scores.
-The strong performance of Deviation patching, as defined in Section 3.3, underscores its ability to better align generated anomalies with true anomaly distributions, making it a robust and effective approach for anomaly generation and detection.
-Table 4. Performance metrics for different patching methods.
-Metric Dataset wo Patch Length-driven Deviation-based F1 MSL 0.535 0.547 0.579 SMAP 0.533 0.571 0.620 Yahoo 0.755 0.741 0.811 AUPR MSL 0.466 0.488 0.529 SMAP 0.425 0.429 0.467 Yahoo 0.776 0.766 0.794
-Figure 5. F1 and AUPR for different patching methods.
-4.8.2. Effectiveness of the Prior Variance
-Figure 6 analyzes the impact of prior variance on anomaly detection performance, showing consistent trends across datasets and indicating that the model is not highly sensitive to this parameter.
-For the best F1 scores, prior variances of 0.5 achieve optimal performance across all datasets (MSL: 0.579, SMAP: 0.62, Yahoo: 0.811), indicating a spot where the model balances bias and variance effectively.
-However, as the prior variance increases beyond 0.5, performance declines, suggesting that overly broad priors introduce excessive flexibility, reducing model specificity.
-Similarly, the AUPR results follow this trend, with 0.5 showing superior means across datasets and narrower standard deviation (std) bars, emphasizing the robustness of this setting.
-The increasing std with higher variances further supports that larger priors may lead to unstable and less reliable performance.
+Table 4: Performance comparison of patching methods
+Dataset wo Patch Length-driven Deviation-based F1 AUPR F1 AUPR F1 AUPR MSL 0.535 0.466 0.547 0.488 0.579 0.529 SMAP 0.533 0.425 0.571 0.429 0.620 0.467 Yahoo 0.755 0.776 0.741 0.766 0.811 0.794
+Figure 5: F1 and AUPR for different patching methods.
+Impact of the Patch Selection.
+Figure 5 presents our analysis of F1 and AUPR for two patching methods—Length-driven and Deviation patching—evaluated across varying patch lengths and thresholds.
+The x-axis indicates the varying portion of patch length and the threshold $\tau$ values for these approaches.
+The best results from this parameter study are compared in Table 4, demonstrating that Deviation patching consistently outperforms other approaches, achieving the highest F1 and AUPR scores across all datasets.
+For example, on SMAP, it improves F1 by 0.044 and AUPR by 0.063 compared to the baseline without patching (wo Patch).
+Length-driven patching, which injects anomalies based on predefined temporal lengths, shows moderate improvements over the baseline in most scenarios but slightly underperforms on Yahoo dataset, where it reduces F1 and AUPR scores.
+In contrast, the strong performance of Deviation patching (Section 3.3) highlights its effectiveness in aligning generated anomalies with true anomaly distributions, making it a more robust and effective approach for anomaly generation and detection.
+Figure 6: F1 and AUPR for different prior variance ($\sigma_{\text{prior}}^{2}$)
+Effectiveness of the Prior Variance.
+Figure 6 analyzes the impact of prior variance on anomaly detection performance, showing consistent trends across datasets and low sensitivity to this parameter.
+A prior variance of 0.5 yields the best F1 scores across all datasets (MSL: 0.579, SMAP: 0.62, Yahoo: 0.811), suggesting an effective balance between bias and variance.
+However, as the prior variance increases beyond 0.5, performance declines, indicating that overly broad priors reduce model specificity.
+Similarly, the AUPR results follow this trend, with 0.5 showing superior means and the smallest standard deviation (std), emphasizing the robustness of this setting.
+The growing std at higher variances further supports that large priors can lead to unstable and less reliable performance.
 Overall, a prior variance of 0.5 emerges as the most effective choice, yielding a balance between accuracy and stability for both F1 and AUPR.
-Figure 6. F1 and AUPR for different prior variance ($\sigma_{\text{prior}}^{2}$)
-4.8.3. Effectiveness of the Min-Max Margin
-Figure 7. Heatmaps of F1 and AUPR scores across datasets illustrating the effectiveness of min-max margins.
-The heatmaps in Figure 7 illustrate the impact of min-max margins ($\delta_{\text{min}},\delta_{\text{max}}$) on performance metrics, such as F1 and AUPR, across MSL, SMAP, and Yahoo datasets.
+Figure 7: Effectiveness of min-max margins.
+Effectiveness of the Min-Max Margin.
+The heatmaps in Figure 7 illustrate the impact of min-max margins ($\delta_{\text{min}},\delta_{\text{max}}$) on F1 across MSL, SMAP, and Yahoo datasets.
 Smaller max margins ($\delta_{\text{max}}=0.2$) and moderate min margins ($\delta_{\text{min}}=0.1$) consistently yield the best results, with peak F1 scores of 0.579, 0.62, and 0.811 for MSL, SMAP, and Yahoo, respectively.
-Similarly, the AUPR values align with this trend.
-These results validate the design rationale in Section 3.2.3, where $\delta_{\text{min}}$ enforces distinctness, and $\delta_{\text{max}}$ maintains realism.
-Larger margins (e.g., $\delta_{\text{max}}=1$) reduce performance, as seen in the lower F1 scores of MSL and Yahoo.
-These findings emphasize the importance of balancing min-max margins.
-5. Conclusion and Future Work
-In this work, we introduced GenIAS, a generative anomaly injection framework designed to enhance time series anomaly detection by creating realistic and diverse anomalies.
+These results validate the design rationale in Section 3.2, where $\delta_{\text{min}}$ enforces distinctness and $\delta_{\text{max}}$ maintains realism.
+Larger margins (e.g., $\delta_{\text{max}}=1$) reduce performance, as seen in lower F1 scores of MSL and Yahoo, emphasizing the importance of balancing margins.
+5 Conclusion and Future Work
+We introduced GenIAS, a generative anomaly injection framework designed to enhance time series anomaly detection by creating realistic and diverse anomalies.
 By perturbing latent representations of a generative model trained on normal data, GenIAS produces anomalies that are similar to real-world anomalies.
-When integrated into a TSAD model, it further enhances anomaly detection performance by adding high-quality synthetic anomalies to the training set.
-For future work, we aim to refine adaptive perturbation techniques and an adaptive patching approach to better align with real-world domain-specific anomalies.
-Additionally, we plan to incorporate correlation between dimensions to guarantee the generation of intermetric anomalies, making the synthetic anomalies more representative of real-world multivariate dependencies.
-These enhancements will improve the realism of generated anomalies by GenIAS and its effectiveness in time series anomaly detection.
+Integrated into a TSAD model, it boosts anomaly detection performance by enriching training data with high-quality synthetic anomalies.
+For future work, we will explore adaptive perturbation and patching strategies, along with modelling inter-dimensional correlations, to further improve synthetic anomaly generation for TSAD.
+Impact Statement
+GenIAS improves TSAD by enabling realistic and controllable synthetic anomaly generation using variational neural models.
+This alleviates the difficulty of training robust algorithms in application domains with few labeled anomalies, particularly for complex multivariate time series with subtle anomalous patterns.
 Appendix A Datasets Descriptions
@@
 You can find the list of symbols and notations, and their definitions in Table 5.
-Table 5. List of Symbols and Notations
+Table 5: List of Symbols and Notations
 Symbol Definition
@@
 $p(\mathbf{X})$ Normal data distribution
-$p(\widetilde{\mathbf{X}})$ Anomalous data distribution
+[…]
 $q(\mathbf{z}|\mathbf{X})$ Latent distribution of normal data
-$p(\widetilde{\mathbf{z}})$ Shifted latent distribution for anomalies
+[…]
 $\mu$ Mean of the latent distribution
@@
 The pseudo-code for GenIAS is provided in Algorithm 1, while the deviation-based patching approach is detailed in Algorithm 2.
+Algorithm 1 End-to-End Training of GenIAS
 Input: TS dataset $\mathcal{D}$, prior standard deviation $\sigma_{\text{prior}}$, loss coefficients $\alpha$, $\beta$, $\gamma$, $\zeta$.
 Output: Trained encoder parameters $\phi$, decoder parameters $\theta$, perturbation scale $\psi$.
-1Initialize encoder $q_{\phi}$, decoder $g_{\theta}$, latent space parameters $\mu$, $\sigma$, and perturbation scale $\psi$;
-2 while training not converged do
-      3 Sample a batch $\{\mathbf{\text{X}}_{i}\}_{i=1}^{B}\subset\mathcal{D}$ of size $B$;
-      4 Encode: Obtain latent parameters $\mu_{i}$, $\sigma_{i}$ by $q_{\phi}(\mathbf{\text{X}}_{i})$;
-      5 Sample latent vectors $\mathbf{\text{z}}_{i}\sim\mathcal{N}({\mu}_{i},\operatorname{diag}({\sigma}_{i}^{2}))$;
-      6 Generate perturbed latent vectors $\widetilde{\mathbf{z}}_{i}$triangleright Eq. (1)
-      7 Decode reconstructed samples: $\hat{\mathbf{X}}_{i}=g_{\theta}(\mathbf{z}_{i})$;
-      8 Decode anomalous samples: $\hat{\mathbf{X}}_{i}\ =g_{\theta}(\widetilde{\mathbf{z}}_{i})$;
-      9 Compute the reconstruction loss $\mathcal{L}_{\text{recon}}$; $\triangleright$ Eq. (2)
-      10 Compute the perturbation loss $\mathcal{L}_{\text{perturb}}$; $\triangleright$ Eq. (3)
-      11 Compute the zero-perturb loss $\mathcal{L}_{\text{zero-perturb}}$; $\triangleright$ Eq. (4)
-      12 Compute the enhanced KL loss $\mathcal{L}_{\text{en-KL}}$; $\triangleright$ Eq. (5)
-      13 $\mathcal{L}_{\text{total}}=\alpha\cdot\mathcal{L}_{\text{recon}}+\beta\cdot\mathcal{L}_{\text{perturb}}+\gamma\cdot\mathcal{L}_{\text{zero-perturb}}+\zeta\cdot\mathcal{L}_{\text{en-KL}}$;
-      14 Update the model parameters $\phi$, $\theta$ and $\psi$  by minimizing $\mathcal{L}_{\text{total}}$;
-15return $\phi$, $\theta$, $\psi$;
-Algorithm 1 End-to-End Training of GenIAS
+Initialize encoder $q_{\phi}$, decoder $g_{\theta}$, latent space parameters $\mu$, $\sigma$, and perturbation scale $\psi$
+while training not converged do
+ 14131211109876543 Sample a batch $\{\mathbf{\text{X}}_{i}\}_{i=1}^{B}\subset\mathcal{D}$ of size $B$  Encode: Obtain latent parameters $\mu_{i}$, $\sigma_{i}$ by $q_{\phi}(\mathbf{\text{X}}_{i})$  Sample latent vectors $\mathbf{\text{z}}_{i}\sim\mathcal{N}({\mu}_{i},\operatorname{diag}({\sigma}_{i}^{2}))$  Generate perturbed latent vectors $\widetilde{\mathbf{z}}_{i}$ $\triangleright$ Eq.
+(1)
+Decode reconstructed samples: $\hat{\mathbf{X}}_{i}=g_{\theta}(\mathbf{z}_{i})$  Decode anomalous samples: $\hat{\mathbf{X}}_{i}\ =g_{\theta}(\widetilde{\mathbf{z}}_{i})$  Compute the reconstruction loss $\mathcal{L}_{\text{recon}}$; $\triangleright$ Eq.
+(5)
+Compute the perturbation loss $\mathcal{L}_{\text{perturb}}$; $\triangleright$ Eq. (3)
+Compute the zero-perturb loss $\mathcal{L}_{\text{zero-perturb}}$; $\triangleright$ Eq. (4)
+Compute the compact KL loss $\mathcal{L}_{\text{comp-KL}}$; $\triangleright$ Eq. (2)
+$\mathcal{L}_{\text{total}}=\alpha\cdot\mathcal{L}_{\text{recon}}+\beta\cdot\mathcal{L}_{\text{perturb}}+\gamma\cdot\mathcal{L}_{\text{zero-perturb}}+\zeta\cdot\mathcal{L}_{\text{comp-KL}}$;
+Update the model parameters $\phi$, $\theta$ and $\psi$ by minimizing $\mathcal{L}_{\text{total}}$
+15return $\phi$, $\theta$, $\psi$
+Algorithm 2 Patching with GenIAS
 Input: Pretrained GenIAS model’s encoder $q_{\phi}$, decoder $g_{\theta}$ and perturbation scales $\psi$, time series window $\mathbf{X}$, threshold scaling factor $\tau$.
@@
 Phase 1: Anomaly Generation
-1 Load encoder $q_{\phi}$, decoder $g_{\theta}$, and perturbation scales $\psi$.
-2Encode: Obtain latent parameters $\mu$, $\sigma$ by $q_{\phi}(\mathbf{\text{X}})$;
-3Generate perturbed latent vectors $\widetilde{\mathbf{z}}$triangleright Eq. (1)
-4Decode $\widetilde{\mathbf{z}}$ to generate preliminary anomalies: $\tilde{\mathbf{X}}=g_{\theta}(\tilde{\mathbf{z}})$;
+Load encoder $q_{\phi}$, decoder $g_{\theta}$, and perturbation scales $\psi$.
+Encode: Obtain latent parameters $\mu$, $\sigma$ by $q_{\phi}(\mathbf{\text{X}})$
+Generate perturbed latent vectors $\widetilde{\mathbf{z}} \triangleright$ Eq. (1)
+[…]
 Phase 2: Deviation-Based Patching
-5 foreach dimension $d$ in $\mathbf{X}$ do
-      6 $\text{deviation}_{d}=\left|\mathbf{X}_{d}-\tilde{\mathbf{X}}_{d}\right|^{2}$;
-      7 $\text{amplitude}_{d}=\max(\mathbf{X}_{d})-\min(\mathbf{X}_{d})$;
-      8 Generate patched output for dimension $d$:
+foreach dimension $d$ in $\mathbf{X}$ do
+ 232221 $\text{deviation}_{d}=\left|\mathbf{X}_{d}-\tilde{\mathbf{X}}_{d}\right|^{2}$$\text{amplitude}_{d}=\max(\mathbf{X}_{d})-\min(\mathbf{X}_{d})$  Generate patched output for dimension $d$:
 $\tilde{\mathbf{X}}_{\text{patched},d}=\begin{cases}\tilde{\mathbf{X}}_{d}&\text{if deviation}_{d}>\tau\cdot\text{amplitude}_{d},\\ \mathbf{X}_{d}&\text{otherwise}.\end{cases}$
-9return $\tilde{\mathbf{X}}_{\text{patched}}$ as the patched anomalous TS windows.
-Algorithm 2 Patching with GenIAS
+24return $\tilde{\mathbf{X}}_{\text{patched}}$ as the patched anomalous TS windows.
 Appendix D Proofs of Theorem
-D.1. Notation and Setup
+D.1 Notation and Setup
 Assume that the latent encoder produces a Gaussian posterior
-(7) $q_{\phi}(z\mid{\mathbf{X}})=\mathcal{N}\left(\mu,\operatorname{diag}(\sigma^{2})\right),$
+$q_{\phi}(z\mid{\mathbf{X}})=\mathcal{N}\left(\mu,\operatorname{diag}(\sigma^{2})\right),$ (7)
 and that the prior is chosen as
-(8) $p(z)=\mathcal{N}\left(0,\sigma^{2}_{\text{prior}}I\right),$
+$p(z)=\mathcal{N}\left(0,\sigma^{2}_{\text{prior}}I\right),$ (8)
 where $\sigma^{2}_{\text{prior}}$ is a hyperparameter and $\mu$ is a vector $\mu=(\mu_{1},...,\mu_{L})$.
-In the following, we consider the effect of using a prior $\sigma^{2}_{\text{prior}}<1$ in our Enhanced KL Loss (Equation 5).
-D.2. Prerequisites
-Lemma 0 (Compact Latent Space in Enhanced KL Loss).
+In the following, we consider the effect of using a prior $\sigma^{2}_{\text{prior}}<1$ in our Compact KL Loss (Equation 2).
+D.2 Prerequisites
+Lemma D.1 (Compact Latent Space in Compact KL Loss).
 Consider the KL divergence between the estimated posterior and the prior distributions for a latent dimension $j$.
-If the enhanced KL loss -Equation 5- is minimized (or its gradient dominates locally) with a prior variance $\sigma^{2}_{\text{prior}}<1$, then for each latent variable $z_{j}$ the optimal posterior variance satisfies $\sigma^{2}_{j}=\sigma^{2}_{\text{prior}}$.
-Hence, the variance of latent representations for normal samples are forced to be less than 1, i.e., more “compact”.
+If the compact KL loss -Equation 2- is minimized (or its gradient dominates locally) with a prior variance $\sigma^{2}_{\text{prior}}<1$, then for each latent variable $z_{j}$ the optimal posterior variance satisfies $\sigma^{2}_{j}=\sigma^{2}_{\text{prior}}$.
+Hence, the variance of latent representations for normal samples is forced to be less than 1, i.e., more “compact”.
 Proof.
 For a one-dimensional Gaussian posterior $\mathcal{N}(\mu,\sigma^{2})$ and a prior $\mathcal{N}(0,\sigma^{2}_{\text{prior}})$, the KL divergence is
-(9) $D_{\text{KL}}(q(z\mid\mathbf{X})\parallel p(z))=\frac{1}{2}\left(\frac{\sigma^{2}}{\sigma^{2}_{\text{prior}}}+\frac{\mu^{2}}{\sigma^{2}_{\text{prior}}}-1-\log\frac{\sigma^{2}}{\sigma^{2}_{\text{prior}}}\right).$
+$D_{\text{KL}}(q(z\mid\mathbf{X})\parallel p(z))=\frac{1}{2}\left(\frac{\sigma^{2}}{\sigma^{2}_{\text{prior}}}+\frac{\mu^{2}}{\sigma^{2}_{\text{prior}}}-1-\log\frac{\sigma^{2}}{\sigma^{2}_{\text{prior}}}\right).$ (9)
 If we minimize the KL term with respect to $\sigma^{2}$ (ignoring the contribution from the reconstruction term and assuming $\mu$ is fixed), we define a function $f$ that takes the part that depends on $\sigma^{2}$:
-(10) $f(\sigma^{2})=\frac{1}{2}\left(\frac{\sigma^{2}}{\sigma^{2}_{\text{prior}}}-\log\frac{\sigma^{2}}{\sigma^{2}_{\text{prior}}}\right).$
+$f(\sigma^{2})=\frac{1}{2}\left(\frac{\sigma^{2}}{\sigma^{2}_{\text{prior}}}-\log\frac{\sigma^{2}}{\sigma^{2}_{\text{prior}}}\right).$ (10)
 Taking the derivative,
-(11) $\frac{df}{d\sigma^{2}}=\frac{1}{2}\left(\frac{1}{\sigma^{2}_{\text{prior}}}-\frac{1}{\sigma^{2}}\right).$
+$\frac{df}{d\sigma^{2}}=\frac{1}{2}\left(\frac{1}{\sigma^{2}_{\text{prior}}}-\frac{1}{\sigma^{2}}\right).$ (11)
 Setting the derivative to zero yields
-(12) $\frac{1}{\sigma^{2}_{\text{prior}}}-\frac{1}{\sigma^{2}}=0\quad\Longrightarrow\quad\sigma^{2}=\sigma^{2}_{\text{prior}}.$
+$\frac{1}{\sigma^{2}_{\text{prior}}}-\frac{1}{\sigma^{2}}=0\quad\Longrightarrow\quad\sigma^{2}=\sigma^{2}_{\text{prior}}.$ (12)
 Thus, when $\sigma^{2}_{\text{prior}}<1$, the optimal posterior variance (in the sense of minimizing the KL divergence) is less than 1, leading to a more compact latent representation. ∎
-Lemma 0 (Reduced Reconstruction Error for Normal Samples).
+Lemma D.2 (Reduced Reconstruction Error for Normal Samples).
 Assume that the decoder $g_{\theta}(z)$ is trained on normal data.
@@
 For an input $\mathbf{X}$ with latent variable $z\sim q_{\phi}(z\mid\mathbf{X})$ and decoder $g_{\theta}(z)$, the reconstruction error is given by
-(13) $\text{MSE}(\mathbf{X})=\mathbb{E}_{z\sim q_{\phi}(z\mid\mathbf{X})}\left[\|\mathbf{X}-g_{\theta}(z)\|^{2}\right].$
+$\text{MSE}(\mathbf{X})=\mathbb{E}_{z\sim q_{\phi}(z\mid\mathbf{X})}\left[\|\mathbf{X}-g_{\theta}(z)\|^{2}\right].$ (13)
 A first-order Taylor expansion of $g_{\theta}(z)$ around the mean $\mu(\mathbf{X})$ gives
-(14) $g_{\theta}(z)\approx g_{\theta}(\mu(\mathbf{X}))+J_{g_{\theta}}(\mu(\mathbf{X}))(z-\mu(\mathbf{X})),$
+$g_{\theta}(z)\approx g_{\theta}(\mu(\mathbf{X}))+J_{g_{\theta}}(\mu(\mathbf{X}))(z-\mu(\mathbf{X})),$ (14)
 where $J_{g_{\theta}}$ is the Jacobian of $g_{\theta}$ at $\mu(\mathbf{X})$.
 We substitute Equation 14 in Equation 13.
-While $\mu(\mathbf{X})$ stays the same in both Enhanced KL and KL cases, $z-\mu(\mathbf{X})$ varies in these two cases.
-Hence we focus on the variable term in this substitution, i.e., $\mathbb{E}\left[\|J_{g_{\theta}}(z-\mu(\mathbf{X}))\|^{2}\right]$.
+While $\mu(\mathbf{X})$ stays the same in both Compact KL and KL cases, $z-\mu(\mathbf{X})$ varies in these two cases.
+Hence, we focus on the variable term in this substitution, i.e., $\mathbb{E}\left[\|J_{g_{\theta}}(z-\mu(\mathbf{X}))\|^{2}\right]$.
 In a multi-dimensional latent space, if the posterior covariance is $\Sigma(\mathbf{X})=\operatorname{diag}(\sigma_{1}^{2}(\mathbf{X}),\dots,\sigma_{d}^{2}(\mathbf{X}))$, then the second-order contribution to the reconstruction error is approximately
-(15) $\mathbb{E}\left[\|J_{g_{\theta}}(z-\mu(\mathbf{X}))\|^{2}\right]=\operatorname{trace}\left(J_{g_{\theta}}^{T}J_{g_{\theta}}\Sigma(\mathbf{X})\right).$
-When the enhanced KL loss forces each $\sigma_{j}^{2}(\mathbf{X})$ to be small (namely $\sigma^{2}_{\text{prior}}<1$), the contribution of the variability term
-(16) $\operatorname{trace}\left(J_{g_{\theta}}^{T}J_{g_{\theta}}\Sigma(\mathbf{X})\right)$
+$\mathbb{E}\left[\|J_{g_{\theta}}(z-\mu(\mathbf{X}))\|^{2}\right]=\operatorname{trace}\left(J_{g_{\theta}}^{T}J_{g_{\theta}}\Sigma(\mathbf{X})\right).$ (15)
+When the compact KL loss forces each $\sigma_{j}^{2}(\mathbf{X})$ to be small (namely $\sigma^{2}_{\text{prior}}<1$), the contribution of the variability term
+$\operatorname{trace}\left(J_{g_{\theta}}^{T}J_{g_{\theta}}\Sigma(\mathbf{X})\right)$ (16)
 is reduced. Hence, the overall reconstruction error decreases. Figure 8 shows the impact of $\sigma^{2}_{\text{prior}}$ on the reconstruction MSE distribution for entity C-1 in MSL dataset. ∎
-D.3. Proof of Theorem 1
+D.3 Proof of Proposition 1
 Statement: Let a VAE be trained on normal data with a latent prior $\mathcal{N}(0,\sigma^{2}_{\text{prior}})$, where $\sigma_{\text{prior}}<1$.
@@
 For generated anomalous samples, assume that the latent posterior is
-(17) $q_{\text{normal}}(z)=\mathcal{N}(\mu_{\text{normal}},\operatorname{diag}(\sigma^{2}_{\text{normal}})),$
-with $\sigma^{2}_{\text{normal}}\approx\sigma^{2}_{\text{prior}}$ (by Lemma 1). For anomalous inputs, suppose the encoder yields an inflated variance so that
-(18) $q_{\text{anom}}(z)=\mathcal{N}(\mu_{\text{anom}},\operatorname{diag}(\psi\sigma^{2}_{\text{normal}})),$
+$q_{\text{normal}}(z)=\mathcal{N}(\mu_{\text{normal}},\operatorname{diag}(\sigma^{2}_{\text{normal}})),$ (17)
+with $\sigma^{2}_{\text{normal}}\approx\sigma^{2}_{\text{prior}}$ (by Lemma 3.1). For anomalous inputs, suppose the encoder yields an inflated variance so that
+$q_{\text{anom}}(z)=\mathcal{N}(\mu_{\text{anom}},\operatorname{diag}(\psi\sigma^{2}_{\text{normal}})),$ (18)
 with $\psi>1$. We assume that the means are similar, i.e. $\mu_{\text{anom}}\approx\mu_{\text{normal}}$, so that the primary difference is in the variances.
 We aim to show that the KL divergence between the latent distributions of normal and anomalous samples strictly increases when the VAE prior variance is compact ($\sigma_{\text{prior}}<1$) compared to when no prior compactness is enforced ($\sigma_{\text{prior}}\geq 1$).
-From the given equation the KL divergence between the anomalous and normal posterior distributions in one dimension is:
-(19) $\begin{split}D_{\text{KL}}\left(\mathcal{N}(\mu,\psi\sigma^{2}_{\text{normal}})\Big{\|}\mathcal{N}(\mu,\sigma^{2}_{\text{prior}})\right)=\\ \frac{1}{2}\left(\frac{\psi\sigma^{2}_{\text{normal}}}{\sigma^{2}_{\text{prior}}}-1-\log\frac{\psi\sigma^{2}_{\text{normal}}}{\sigma^{2}_{\text{prior}}}\right).\end{split}$
+From the given equation, the KL divergence between the anomalous and normal posterior distributions in one dimension is:
+$\begin{split}D_{\text{KL}}\left(\mathcal{N}(\mu,\psi\sigma^{2}_{\text{normal}})\Big\|\mathcal{N}(\mu,\sigma^{2}_{\text{prior}})\right)=\\ \frac{1}{2}\left(\frac{\psi\sigma^{2}_{\text{normal}}}{\sigma^{2}_{\text{prior}}}-1-\log\frac{\psi\sigma^{2}_{\text{normal}}}{\sigma^{2}_{\text{prior}}}\right).\end{split}$ (19)
 We compare two scenarios:
-  • Scenario A (Compact Latent Space): $\sigma_{\text{prior}}<1$, let $\sigma_{\text{prior}}=\alpha$, where $10<\alpha<$ 1.
+  • Scenario A (Compact Latent Space): $\sigma_{\text{prior}}<1$, let $\sigma_{\text{prior}}=\alpha$, where $0<\alpha<1$.
   • Scenario B (Non-Compact Latent Space): $\sigma_{\text{prior}}\geq 1$, specifically, we consider $\sigma_{\text{prior}}=1$.
 For Scenario A ($\sigma_{\text{prior}}=\alpha$):
-(20) $D_{\text{KL},A}=\frac{1}{2}\left(\frac{\psi\sigma^{2}_{\text{normal}}}{\alpha^{2}}-1-\log\frac{\psi\sigma^{2}_{\text{normal}}}{\alpha^{2}}\right).$
+$D_{\text{KL},A}=\frac{1}{2}\left(\frac{\psi\sigma^{2}_{\text{normal}}}{\alpha^{2}}-1-\log\frac{\psi\sigma^{2}_{\text{normal}}}{\alpha^{2}}\right).$ (20)
 For Scenario B ($\sigma_{\text{prior}}=1$):
-(21) $D_{\text{KL},B}=\frac{1}{2}\left(\frac{\psi\sigma^{2}_{\text{normal}}}{1}-1-\log\frac{\psi\sigma^{2}_{\text{normal}}}{1}\right).$
-Since $10<\alpha<$ 1, we observe that the fraction $\frac{\psi\sigma^{2}_{\text{normal}}}{\alpha^{2}}$ in $D_{\text{KL},A}$ is strictly larger than $\frac{\psi\sigma^{2}_{\text{normal}}}{1}$ in $D_{\text{KL},B}$.
+$D_{\text{KL},B}=\frac{1}{2}\left(\frac{\psi\sigma^{2}_{\text{normal}}}{1}-1-\log\frac{\psi\sigma^{2}_{\text{normal}}}{1}\right).$ (21)
+Since $0<\alpha<1$, we observe that the fraction $\frac{\psi\sigma^{2}_{\text{normal}}}{\alpha^{2}}$ in $D_{\text{KL},A}$ is strictly larger than $\frac{\psi\sigma^{2}_{\text{normal}}}{1}$ in $D_{\text{KL},B}$.
 The function $f(x)=x-1-\log x$ is monotonically increasing for $x>1$, meaning:
-(22) $f\left(\frac{\psi\sigma^{2}_{\text{normal}}}{\alpha^{2}}\right)>f\left(\frac{\psi\sigma^{2}_{\text{normal}}}{1}\right).$
+$f\left(\frac{\psi\sigma^{2}_{\text{normal}}}{\alpha^{2}}\right)>f\left(\frac{\psi\sigma^{2}_{\text{normal}}}{1}\right).$ (22)
 Thus, we obtain:
-(23) $D_{\text{KL},A}>D_{\text{KL},B}.$
+$D_{\text{KL},A}>D_{\text{KL},B}.$ (23)
 Since $D_{\text{KL},A}>D_{\text{KL},B}$, we conclude that the KL divergence between normal and anomalous latent distributions is strictly larger when the VAE is trained with a compact latent space ($\sigma_{\text{prior}}<1$).
-This means that a compact latent space amplifies the separation between normal and anomalous samples under perturbation.
+This means that a compact latent space amplifies the separation between normal and anomalous samples under perturbation. ∎
 Appendix E Details of Evaluation Metrics
-To support a consistent and meaningful evaluation of generated anomalies, we utilize a Deep SVDD (Ruff et al., 2018) representation with a fixed 128-dimensional embedding for all models.
-This representation provides a structured feature space.
-Additionally, we evaluate models using test datasets that contain labeled anomalies, allowing for a more reliable assessment of anomaly generation quality.
-Let $\Omega$ denote the DeepSVDD model.
-The representation of real anomalous samples in the test dataset is given by $V_{\text{real}}=\Omega(\text{X}_{\text{real anomalies}})$, while the representation of generated anomalous samples, produced by the anomaly generator model from the train dataset, is given by $V_{\text{gen}}=\Omega(\tilde{{\text{X}}})$.
-E.1. Anomalous Representation Proximity (ARP)
+[…]
+E.1 Anomalous Representation Proximity (ARP)
 ARP measures how well the generated anomalies cover the distribution of real anomalies in the representation space.
@@
 Formally, if $V_{\text{real}}$ and $V_{\text{gen}}$ are the sets of real and generated anomaly representations, the ARP is defined as:
-(24) $\text{ARP}=\frac{1}{1+\frac{1}{|V_{\text{real}}|}\sum_{\mathbf{v}^{r}\in V_{\text{real}}}\min_{\mathbf{v}^{g}\in V_{\text{gen}}}d(\mathbf{v}^{r},\mathbf{v}^{g})}.$
+$\text{ARP}=\frac{1}{1+\frac{1}{|V_{\text{real}}|}\sum_{\mathbf{v}^{r}\in V_{\text{real}}}\min_{\mathbf{v}^{g}\in V_{\text{gen}}}d(\mathbf{v}^{r},\mathbf{v}^{g})},$ (24)
 where $d(\cdot,\cdot)$ is the Euclidean distance metric. A higher ARP indicates that the generated anomalies are more similar to real anomalies in the representation space, meaning better realism.
-E.2. Entropy-Based Diversity Index (EDI)
+E.2 Entropy-Based Diversity Index (EDI)
 EDI quantifies how well a model generates a diverse set of anomalies by evaluating their distribution across the representation space.
 First, we collect the representations of all generated anomalies from all models into a global set:
-(25) $V_{\text{all}}=\bigcup_{m=1}^{M}V_{\text{gen}}^{(m)},$
+$V_{\text{all}}=\bigcup_{m=1}^{M}V_{\text{gen}}^{(m)},$ (25)
 where $V_{\text{gen}}^{(m)}=\{\mathbf{v}_{1}^{(m)},\mathbf{v}_{2}^{(m)},\dots,\mathbf{v}_{N_{m}}^{(m)}\}$ represents the set of generated anomalies for model $m$, and $M$ is the number of models being evaluated.
@@
 For a specific model $m$, the proportion of its generated anomalies in region $R_{i}$ is defined as:
-(26) $p_{i}^{(m)}=\frac{|\{\mathbf{v}_{j}^{(m)}\in V_{\text{gen}}^{(m)}\mid\mathbf{v}_{j}^{(m)}\in R_{i}\}|}{N_{m}},$
-where $N_{m}=|V_{\text{gen}}^{(m)}|$ is the total number of generated anomalies for model $m$, and ||$\cdot$|| ⋅ | denotes the cardinality of a set.
+$p_{i}^{(m)}=\frac{|\{\mathbf{v}_{j}^{(m)}\in V_{\text{gen}}^{(m)}\mid\mathbf{v}_{j}^{(m)}\in R_{i}\}|}{N_{m}},$ (26)
+where $N_{m}=|V_{\text{gen}}^{(m)}|$ is the total number of generated anomalies for model $m$, and $|\cdot|$ denotes the cardinality of a set.
 The EDI for model $m$ is then computed using Shannon Entropy (Shannon, 1948) as:
-(27) $\text{EDI}^{(m)}=-\sum_{i=1}^{K}p_{i}^{(m)}\log p_{i}^{(m)}.$
+$\text{EDI}^{(m)}=-\sum_{i=1}^{K}p_{i}^{(m)}\log p_{i}^{(m)}.$ (27)
 A higher EDI indicates that a model generates anomalies that are well-distributed across the space, reflecting greater diversity rather than being concentrated in a few regions.
-E.3. TSAD Evaluation Metrics
-We use Best F1 score to measures the performance of TSAD models at the optimal threshold.
+E.3 TSAD Evaluation Metrics
+We use the Best F1 score to measure the performance of TSAD models at the optimal threshold.
 AUPR and AUROC are holistic measures that assess model performance across various thresholds.
@@
 All evaluations were performed on a system with an A40 GPU, 13 CPU cores, and 250 GB of RAM. We use the same hyperparameters for GenIAS across UTS and MTS datasets, as detailed in Table 6.
-Table 6. hyperparameter configurations used in GenIAS.
+Table 6: hyperparameter configurations used in GenIAS.
 Parameter Value
@@
 $\sigma_{\text{prior}}$ 0.5
-lr initial=$10^{-4}$, with scheduled reduction
+[…]
 epochs max=1000
@@
 Appendix G Additional Results
-G.1. Visualizing Enhanced KL Loss Effects
+G.1 Visualizing Compact KL Loss Effects
 Separation of reconstruction error in the input space is illustrated in Figure 8.
-Figure 8. Separation of reconstruction error (anomaly score) in the input space with a prior of 0.5 (left) and without a prior (right) on entity C-1 from the MSL dataset. The y-axis is logarithmic.
-G.2. TSAD additional results
-G.2.1. Critical Difference diagrams
-Statistical significance in critical difference digrams was determined using a Friedman test followed by a post-hoc Wilcoxon signed-rank test with Holm correction (Demšar, 2006).
+Figure 8: Separation of reconstruction error (anomaly score) in the input space with a prior of 0.5 (left) and without a prior (right) on entity C-1 from the MSL dataset. The y-axis is logarithmic.
+G.2 TSAD additional results
+G.2.1 Critical Difference diagrams
+Statistical significance in critical difference diagrams was determined using a Friedman test followed by a post-hoc Wilcoxon signed-rank test with Holm correction (Demšar, 2006).
 The critical difference diagrams in Figure 9 show that GenIAS consistently ranks first across both multivariate (MTS) and univariate (UTS) time series settings for AUPR and F1 Score.
 In the MTS setting, GenIAS demonstrates a statistically significant improvement over other baselines, indicating its superior performance in capturing anomalies.
-Figure 9. Critical difference diagrams for F1 and AUPR scores: (a, b) represent multivariate models, while (c, d) showcase univariate models.
-G.2.2. TSAD performance in the other metrics.
+Figure 9: Critical difference diagrams for F1 and AUPR: (a, b) represent multivariate models, while (c, d) showcase univariate models.
+G.2.2 TSAD performance in the other metrics.
 In Table 7, we present the TSAD performance of GenIAS and the baseline methods in terms of three additional metrics: Affiliation F1 (Aff), PATE, and AUROC.
@@
 In terms of AUROC, GenIAS achieves the second-best overall performance and best overall performance among neural network-based methods, while Isolation Forest, a well-known legacy shallow AD method, yields surprisingly good performance.
-Table 7.
-TSAD performance of GenIAS and baseline methods in terms of Aff, PATE and AUROC, with the best model in bold and the second-best model underlined.
+Table 7: TSAD performance of GenIAS and baseline methods in terms of Aff, PATE, and AUROC, with the best model in bold and the second-best model is underlined.
 U and M denote results unavailable for models limited to UTS or MTS, respectively.
 Multivariate Univariate Metric Cat.
-Methods MSL SMAP SMD SWAT GECCO SWAN Avg-M Yahoo KPI UCR Avg-U Avg Aff Classic OCSVM (1999) 0.742 0.761 0.733 0.197 0.425 0.711 0.595 0.923 0.761 0.710 0.798 0.663 LOF (2000) 0.775 0.762 0.666 0.613 0.227 0.745 0.631 0.906 0.674 0.718 0.766 0.676 IForest (2008) 0.719 0.728 0.699 0.535 0.259 0.824 0.627 0.907 0.760 0.686 0.784 0.680 PU LSTM-VAE (2018) 0.824 0.782 0.630 0.324 0.312 0.726 0.600 M M M M M US-WI Donut (2018) U U U U U U U 0.777 0.682 0.685 0.715 - Omni (2019) 0.680 0.779 0.706 0.156 0.230 0.608 0.527 M M M M M THOC (2020) 0.778 0.781 0.677 0.679 0.321 0.713 0.658 0.796 0.676 0.642 0.705 0.674 MTAD-GAT(2020) 0.833 0.848 0.731 0.355 0.308 0.745 0.637 M M M M M AnomalyTran (2021) 0.774 0.766 0.664 0.482 0.551 0.537 0.629 M M M M M TranAD (2022) 0.771 0.751 0.719 0.606 0.233 0.661 0.624 0.917 0.779 0.698 0.798 0.682 TS2Vec (2022) 0.694 0.695 0.750 0.700 0.520 0.728 0.681 0.867 0.818 0.671 0.785 0.716 TimesNet (2023) 0.769 0.710 0.795 0.693 0.829 0.710 0.751 0.917 0.844 0.712 0.824 0.775 Dcdetector (2023) 0.706 0.680 0.528 0.693 0.671 0.710 0.665 0.737 0.666 0.671 0.691 0.674 US-I NCAD (2022) 0.685 0.664 0.732 0.693 0.887 0.710 0.729 0.645 0.753 0.747 0.715 0.724 CutAddPaste (2024) 0.762 0.753 0.693 0.764 0.659 0.875 0.751 0.832 0.818 0.820 0.823 0.775 COUTA (2024) 0.785 0.762 0.764 0.645 0.592 0.731 0.713 0.906 0.851 0.730 0.829 0.752 CARLA (2025) 0.816 0.808 0.733 0.592 0.756 0.932 0.773 0.874 0.701 0.805 0.793 0.780 Ours GenIAS 0.851 0.856 0.731 0.596 0.763 0.935 0.789 0.921 0.788 0.814 0.841 0.806 PATE Classic OCSVM (1999) 0.267 0.270 0.363 0.698 0.191 0.704 0.416 0.780 0.343 0.033 0.385 0.405 LOF (2000) 0.242 0.251 0.319 0.187 0.042 0.706 0.291 0.690 0.159 0.056 0.302 0.295 IForest (2008) 0.195 0.197 0.306 0.743 0.191 0.799 0.405 0.735 0.325 0.027 0.362 0.391 PU LSTM-VAE (2018) 0.318 0.278 0.436 0.693 0.061 0.753 0.423 M M M M M US-WI Donut (2018) U U U U U U U 0.358 0.120 0.018 0.165 - Omni (2019) 0.143 0.162 0.422 0.702 0.076 0.709 0.369 M M M M M THOC (2020) 0.307 0.218 0.147 0.508 0.104 0.563 0.308 0.488 0.317 0.016 0.274 0.296 MTAD-GAT(2020) 0.391 0.371 0.445 0.115 0.247 0.707 0.379 M M M M M AnomalyTran (2021) 0.267 0.283 0.332 0.683 0.112 0.685 0.394 M M M M M TranAD (2022) 0.238 0.279 0.458 0.192 0.141 0.688 0.333 0.777 0.408 0.025 0.403 0.356 TS2Vec (2022) 0.143 0.150 0.137 0.147 0.055 0.530 0.194 0.636 0.247 0.144 0.342 0.243 TimesNet (2023) 0.282 0.207 0.452 0.083 0.240 0.636 0.317 0.760 0.355 0.042 0.386 0.340 Dcdetector (2023) 0.176 0.135 0.097 0.124 0.025 0.455 0.169 0.184 0.052 0.025 0.087 0.141 US-I NCAD (2022) 0.157 0.161 0.129 0.107 0.272 0.556 0.230 0.145 0.173 0.130 0.149 0.203 CutAddPaste (2024) 0.269 0.347 0.200 0.613 0.252 0.842 0.421 0.237 0.741 0.262 0.413 0.418 COUTA (2024) 0.270 0.244 0.447 0.159 0.285 0.756 0.360 0.684 0.335 0.054 0.358 0.359 CARLA (2025) 0.493 0.413 0.488 0.655 0.193 0.911 0.526 0.695 0.378 0.301 0.458 0.503 Ours GenIAS 0.571 0.516 0.553 0.718 0.274 0.941 0.596 0.838 0.456 0.331 0.542 0.578 AUC-ROC Classic OCSVM (1999) 0.605 0.634 0.762 0.799 0.446 0.714 0.660 0.918 0.728 0.567 0.738 0.686 LOF (2000) 0.610 0.584 0.708 0.660 0.370 0.789 0.620 0.890 0.604 0.575 0.690 0.643 IForest (2008) 0.571 0.579 0.768 0.842 0.927 0.862 0.758 0.915 0.749 0.570 0.745 0.754 PU LSTM-VAE (2018) 0.618 0.650 0.771 0.806 0.506 0.763 0.686 M M M M M US-WI Donut (2018) U U U U U U U 0.676 0.640 0.498 0.605 - Omni (2019) 0.387 0.570 0.794 0.811 0.430 0.696 0.615 M M M M M THOC (2020) 0.654 0.604 0.678 0.318 0.395 0.578 0.538 0.873 0.799 0.558 0.743 0.606 MTAD-GAT(2020) 0.674 0.641 0.812 0.822 0.590 0.749 0.715 M M M M M AnomalyTran (2021) 0.641 0.653 0.764 0.809 0.698 0.678 0.707 M M M M M TranAD (2022) 0.632 0.645 0.808 0.677 0.519 0.677 0.660 0.889 0.763 0.515 0.722 0.681 TS2Vec (2022) 0.503 0.517 0.496 0.548 0.501 0.573 0.523 0.698 0.567 0.501 0.589 0.545 TimesNet (2023) 0.664 0.578 0.837 0.243 0.924 0.598 0.641 0.892 0.750 0.597 0.746 0.676 Dcdetector (2023) 0.535 0.519 0.579 0.504 0.504 0.499 0.523 0.603 0.504 0.509 0.539 0.528 US-I NCAD (2022) 0.491 0.493 0.572 0.365 0.692 0.592 0.534 0.567 0.671 0.609 0.616 0.561 CutAddPaste (2024) 0.673 0.631 0.631 0.777 0.789 0.464 0.661 0.684 0.810 0.617 0.704 0.675 COUTA (2024) 0.661 0.606 0.811 0.643 0.660 0.802 0.697 0.911 0.787 0.559 0.752 0.716 CARLA (2025) 0.694 0.677 0.657 0.783 0.696 0.510 0.670 0.543 0.604 0.735 0.627 0.655 Ours GenIAS 0.745 0.718 0.633 0.786 0.811 0.536 0.705 0.778 0.691 0.778 0.749 0.720
-(a) KPI - Entity c69a50cf-ee03-3bd7-831e-407d36c7ee91
-(b) SMD - Entity machine-2-3
-(c) UCR - Entity 222_UCR_Anomaly_mit14046longtermecg_56123_91200_91700
-Figure 10.
-Each plot visualises a different time series (TS) injection or generation method using t-SNE.
+Methods MSL SMAP SMD SWAT GECCO SWAN Avg-M Yahoo KPI UCR Avg-U Avg AUPR Classic OCSVM (1999) 0.202 0.256 0.327 0.696 0.163 0.608 0.375 0.703 0.300 0.028 0.344 0.365 LOF (2000) 0.217 0.248 0.277 0.185 0.031 0.624 0.264 0.615 0.109 0.043 0.256 0.261 IForest (2008) 0.182 0.190 0.278 0.739 0.162 0.739 0.382 0.614 0.288 0.021 0.308 0.357 PU LSTM-VAE (2018) 0.285 0.258 0.395 0.685 0.045 0.674 0.390 M M M M M US-WI Donut (2018) U U U U U U U 0.264 0.078 0.015 0.119 U Omni (2019) 0.149 0.115 0.365 0.713 0.034 0.587 0.327 M M M M M THOC (2020) 0.240 0.195 0.107 0.537 0.076 0.432 0.265 0.349 0.229 0.014 0.197 0.242 MTAD-GAT(2020) 0.335 0.339 0.401 0.095 0.200 0.613 0.331 M M M M M AnomalyTran (2021) 0.236 0.264 0.273 0.680 0.059 0.568 0.347 M M M M M TranAD (2022) 0.273 0.287 0.412 0.192 0.119 0.586 0.312 0.691 0.285 0.022 0.333 0.319 TS2Vec (2022) 0.132 0.148 0.113 0.136 0.040 0.408 0.163 0.491 0.221 0.009 0.240 0.189 TimesNet (2023) 0.283 0.208 0.385 0.082 0.189 0.508 0.276 0.671 0.237 0.032 0.313 0.288 Dcdetector (2023) 0.129 0.124 0.043 0.126 0.011 0.326 0.127 0.041 0.018 0.009 0.023 0.092 US-I NCAD (2022) 0.146 0.156 0.096 0.106 0.135 0.407 0.174 0.067 0.089 0.103 0.086 0.145 CutAddPaste (2024) 0.242 0.322 0.153 0.582 0.159 0.808 0.378 0.116 0.602 0.188 0.302 0.352 COUTA (2024) 0.247 0.248 0.400 0.159 0.264 0.685 0.334 0.555 0.282 0.049 0.295 0.321 CARLA (2025) 0.501 0.448 0.507 0.681 0.201 0.814 0.525 0.627 0.299 0.247 0.391 0.481 Ours GenIAS+CARLA 0.529 0.467 0.512 0.704 0.236 0.882 0.555 0.794 0.357 0.280 0.477 0.529 Aff Classic OCSVM (1999) 0.742 0.761 0.733 0.197 0.425 0.711 0.595 0.923 0.761 0.710 0.798 0.663 LOF (2000) 0.775 0.762 0.666 0.613 0.227 0.745 0.631 0.906 0.674 0.718 0.766 0.676 IForest (2008) 0.719 0.728 0.699 0.535 0.259 0.824 0.627 0.907 0.760 0.686 0.784 0.680 PU LSTM-VAE (2018) 0.824 0.782 0.630 0.324 0.312 0.726 0.600 M M M M M US-WI Donut (2018) U U U U U U U 0.777 0.682 0.685 0.715 - Omni (2019) 0.680 0.779 0.706 0.156 0.230 0.608 0.527 M M M M M THOC (2020) 0.778 0.781 0.677 0.679 0.321 0.713 0.658 0.796 0.676 0.642 0.705 0.674 MTAD-GAT(2020) 0.833 0.848 0.731 0.355 0.308 0.745 0.637 M M M M M AnomalyTran (2021) 0.774 0.766 0.664 0.482 0.551 0.537 0.629 M M M M M TranAD (2022) 0.771 0.751 0.719 0.606 0.233 0.661 0.624 0.917 0.779 0.698 0.798 0.682 TS2Vec (2022) 0.694 0.695 0.750 0.700 0.520 0.728 0.681 0.867 0.818 0.671 0.785 0.716 TimesNet (2023) 0.769 0.710 0.795 0.693 0.829 0.710 0.751 0.917 0.844 0.712 0.824 0.775 Dcdetector (2023) 0.706 0.680 0.528 0.693 0.671 0.710 0.665 0.737 0.666 0.671 0.691 0.674 US-I NCAD (2022) 0.685 0.664 0.732 0.693 0.887 0.710 0.729 0.645 0.753 0.747 0.715 0.724 CutAddPaste (2024) 0.762 0.753 0.693 0.764 0.659 0.875 0.751 0.832 0.818 0.820 0.823 0.775 COUTA (2024) 0.785 0.762 0.764 0.645 0.592 0.731 0.713 0.906 0.851 0.730 0.829 0.752 CARLA (2025) 0.816 0.808 0.733 0.592 0.756 0.932 0.773 0.874 0.701 0.805 0.793 0.780 Ours GenIAS+CARLA 0.851 0.856 0.731 0.596 0.763 0.935 0.789 0.921 0.788 0.814 0.841 0.806 PATE Classic OCSVM (1999) 0.267 0.270 0.363 0.698 0.191 0.704 0.416 0.780 0.343 0.033 0.385 0.405 LOF (2000) 0.242 0.251 0.319 0.187 0.042 0.706 0.291 0.690 0.159 0.056 0.302 0.295 IForest (2008) 0.195 0.197 0.306 0.743 0.191 0.799 0.405 0.735 0.325 0.027 0.362 0.391 PU LSTM-VAE (2018) 0.318 0.278 0.436 0.693 0.061 0.753 0.423 M M M M M US-WI Donut (2018) U U U U U U U 0.358 0.120 0.018 0.165 - Omni (2019) 0.143 0.162 0.422 0.702 0.076 0.709 0.369 M M M M M THOC (2020) 0.307 0.218 0.147 0.508 0.104 0.563 0.308 0.488 0.317 0.016 0.274 0.296 MTAD-GAT(2020) 0.391 0.371 0.445 0.115 0.247 0.707 0.379 M M M M M AnomalyTran (2021) 0.267 0.283 0.332 0.683 0.112 0.685 0.394 M M M M M TranAD (2022) 0.238 0.279 0.458 0.192 0.141 0.688 0.333 0.777 0.408 0.025 0.403 0.356 TS2Vec (2022) 0.143 0.150 0.137 0.147 0.055 0.530 0.194 0.636 0.247 0.144 0.342 0.243 TimesNet (2023) 0.282 0.207 0.452 0.083 0.240 0.636 0.317 0.760 0.355 0.042 0.386 0.340 Dcdetector (2023) 0.176 0.135 0.097 0.124 0.025 0.455 0.169 0.184 0.052 0.025 0.087 0.141 US-I NCAD (2022) 0.157 0.161 0.129 0.107 0.272 0.556 0.230 0.145 0.173 0.130 0.149 0.203 CutAddPaste (2024) 0.269 0.347 0.200 0.613 0.252 0.842 0.421 0.237 0.741 0.262 0.413 0.418 COUTA (2024) 0.270 0.244 0.447 0.159 0.285 0.756 0.360 0.684 0.335 0.054 0.358 0.359 CARLA (2025) 0.493 0.413 0.488 0.655 0.193 0.911 0.526 0.695 0.378 0.301 0.458 0.503 Ours GenIAS+CARLA 0.571 0.516 0.553 0.718 0.274 0.941 0.596 0.838 0.456 0.331 0.542 0.578 AUC-ROC Classic OCSVM (1999) 0.605 0.634 0.762 0.799 0.446 0.714 0.660 0.918 0.728 0.567 0.738 0.686 LOF (2000) 0.610 0.584 0.708 0.660 0.370 0.789 0.620 0.890 0.604 0.575 0.690 0.643 IForest (2008) 0.571 0.579 0.768 0.842 0.927 0.862 0.758 0.915 0.749 0.570 0.745 0.754 PU LSTM-VAE (2018) 0.618 0.650 0.771 0.806 0.506 0.763 0.686 M M M M M US-WI Donut (2018) U U U U U U U 0.676 0.640 0.498 0.605 - Omni (2019) 0.387 0.570 0.794 0.811 0.430 0.696 0.615 M M M M M THOC (2020) 0.654 0.604 0.678 0.318 0.395 0.578 0.538 0.873 0.799 0.558 0.743 0.606 MTAD-GAT(2020) 0.674 0.641 0.812 0.822 0.590 0.749 0.715 M M M M M AnomalyTran (2021) 0.641 0.653 0.764 0.809 0.698 0.678 0.707 M M M M M TranAD (2022) 0.632 0.645 0.808 0.677 0.519 0.677 0.660 0.889 0.763 0.515 0.722 0.681 TS2Vec (2022) 0.503 0.517 0.496 0.548 0.501 0.573 0.523 0.698 0.567 0.501 0.589 0.545 TimesNet (2023) 0.664 0.578 0.837 0.243 0.924 0.598 0.641 0.892 0.750 0.597 0.746 0.676 Dcdetector (2023) 0.535 0.519 0.579 0.504 0.504 0.499 0.523 0.603 0.504 0.509 0.539 0.528 US-I NCAD (2022) 0.491 0.493 0.572 0.365 0.692 0.592 0.534 0.567 0.671 0.609 0.616 0.561 CutAddPaste (2024) 0.673 0.631 0.631 0.777 0.789 0.464 0.661 0.684 0.810 0.617 0.704 0.675 COUTA (2024) 0.661 0.606 0.811 0.643 0.660 0.802 0.697 0.911 0.787 0.559 0.752 0.716 CARLA (2025) 0.694 0.677 0.657 0.783 0.696 0.510 0.670 0.543 0.604 0.735 0.627 0.655 Ours GenIAS+CARLA 0.745 0.718 0.633 0.786 0.811 0.536 0.705 0.778 0.691 0.778 0.749 0.720
+(a) SMAP - Entity E5
+(b) KPI - Entity c69a50cf-ee03-3bd7-831e-407d36c7ee91
+(c) SMD - Entity machine-2-3
+(d) UCR - Entity 222_UCR_Anomaly_mit14046longtermecg_56123_91200_91700
+Figure 10: Each plot visualizes a different time series (TS) injection or generation method using t-SNE.
 Blue markers represent the original anomalous TS, while orange markers represent the injected or generated anomalous TS.
-G.2.3. Additional Visualizations
+G.2.3 Additional Visualizations
 To further illustrate the realistic anomaly generation of GenIAS, we present three additional visualizations comparing generated and real anomalies from both GenIAS and the US-i baselines in Figure 10.